### v1

# Designing Adaptive User Interfaces for mHealth applications targeting chronic disease: A User-Centric Approach

Abstract.
mHealth interventions show significant potential to help in the self-management of chronic diseases, but their under use remains a problem. Considering the substantial diversity among individuals dealing with chronic diseases, tailored strategies are essential. Adaptive User Interfaces (AUIs) may help address the diverse and evolving needs of this demographic. To investigate this approach, we developed an AUI prototype informed by existing literature findings. We then used this prototype as the basis for focus group discussions and interview studies with 22 participants managing various chronic diseases, and follow-up surveys of all participants. Through these investigations, we pinpointed key challenges related to the use of AUIs, strategies to improve adaptation design, and potential trade-offs between these challenges and strategies. Concurrently, a quantitative survey was conducted to extract preferences for AUIs in chronic disease-related applications with 90 further participants. This uncovered participants’ preferences for various adaptations, data types, collection methods, and involvement levels. Finally, we synthesised these insights and categories, aligning them with existing guidelines and design considerations for mHealth app adaptation design. This resulted in nine guidelines that we refined by a final feedback survey conducted with 20 participants.
1. Introduction
Chronic diseases, including conditions such as asthma, heart disease, and diabetes, present formidable challenges to health care systems around the world (WHO, 2022). The management of these long-term health conditions transcends simple medical treatment, with an increasing emphasis on empowering patients to actively engage in self-management practices (WHO, 2022). The use of mHealth technologies has emerged as a promising avenue for promoting self-management by strengthening medication adherence and facilitating self-tracking capabilities (Hamine et al., 2015). However, despite the potential of these technologies to improve health outcomes, research indicates that a significant proportion of people who benefit from mHealth solutions do not fully take advantage of them (Han et al., 2010). In order to increase the adoption of mHealth applications, it is essential to develop more user-friendly systems capable of meeting the varied needs of users (Han et al., 2010). However, several challenges need to be addressed to achieve this objective. Firstly, chronic diseases are inherently heterogeneous, impacting individuals in various ways, including triggers, symptoms, and severity (Harvey et al., 2012; Audulv, 2013). Second, it is crucial to consider the development stages of chronic diseases when designing mHealth technology, as chronic diseases can cooccur with other medical or psychological disorders, further adding to the complexity of self-management of chronic diseases (di San Marco et al., 2019; Audulv, 2013). Third, chronic diseases generally persist over an individual’s lifetime (WHO, 2022; Harvey et al., 2012). Therefore, mHealth technologies are needed to keep users engaged and motivated in the long run. This makes chronic disease management beyond the narrow definition of disease management.
Adaptive User Interfaces (AUIs) have often been recommended to accommodate the considerable variability between patients by tailoring the user interface (UI) to align with individual needs, objectives and usage contexts in mHealth technologies (Norcio and Stanley, 1989). Despite increasing interest in employing AUIs within chronic disease related applications (Setiawan et al., 2019; Talboom-Kamp et al., 2018), they often overlook diverse user characteristics and interactions (Grua et al., 2020; McLean et al., 2011). Furthermore, there remains a dearth of information regarding the development processes for AUIs during the early design phases (Wang et al., 2023b). Current AUI evaluation methods focus mainly on evaluating the overall effectiveness of the application, making it difficult to determine the utility of AUIs and user preferences (Wang et al., 2023b). Furthermore, there is limited information on how people with chronic diseases use AUIs, potentially resulting in the underutilisation of adaptation benefits (Kornfield et al., 2020). Addressing these gaps is crucial to optimising AUI design and improving user engagement in chronic disease management applications.
In this work, we examine the perspectives of users on the usage of AUIs in applications related to chronic diseases. First, we develop a prototype that focuses on AUIs in chronic disease-related applications. This prototype served as the foundation for our qualitative investigation into how individuals experience AUIs within the context of chronic diseases, through interviews and focus group studies. Additionally, all participants also completed a post-session survey to provide additional information on adaptations. During the same time, we conducted a quantitative survey to gather user preferences regarding various aspects of the adaptation process. By collecting and analysing qualitative and quantitative data and aligning with the existing literature, we established guidelines for designing AUIs for chronic disease-related applications. These guidelines were refined based on feedback obtained from a subsequent survey, resulting in the finalisation of nine guidelines.
We presented the initial findings of our user study at the 2024 International Conference on Software Engineering (ICSE) (Wang et al., 2024). In this paper, we significantly expand on our previous work by: (i) conducting a detailed and extended analysis of the emerging concepts identified in the user study, incorporating post-session survey data, and mapping the analysis between categories obtained from the study; (ii) enhancing our survey findings by including all collected data and conducting further analysis and interpretation; (iii) developing a set of new guidelines for designing AUIs in four stages; and (iv) refining our guidelines through an additional round of feedback collection. This work offers five key research contributions:
An investigation of the needs and preferences of users with chronic diseases for their mHealth app AUIs, expanding the adaptation literature by uncovering challenges in AUI usage and suggesting practical strategies for improved AUI design;
Advancement of the discussion on this topic by presenting design trade-offs in AUIs when designing technology for users with chronic disease;
Deeper understanding of how user preferences for different aspects of adaptations are influenced by diiferent demographic factors, including cultural backgrounds, contextual circumstances, Health condition and age.
Development of comprehensive actionable guidelines for researchers, practitioners, and designers for designing AUIs in chronic disease domain; and
Initial evaluation of these guidelines with a cohort of 20 users and an iterative refinement of the guidelines.
The rest of the paper is organised as follows. Section 2 provides a summary of our study’s motivation and summarises key related work. In Section 3, we present our research methodology. The results of qualitative and quantitative data analysis are presented in Sections 4 and 5, respectively. Section 6 offers an in-depth description of the proposed guidelines. We present the results from the feedback survey for the guidelines in Section 6.2. Finally, section 7 presents the limitations and threats to the validity of the research and Section 8 concludes this paper.
2. Motivation and Related Work
Adaptive User Interfaces: Due to the diversity of users and usage contexts, a User Interface (UI) developed for a fixed context of use may not be sufficient. AUIs adapt dynamically to users’ contexts and preferences, showing promise in addressing such issues. McTear (1993) defines an AUI as “a software artefact that improves its ability to interact with a user by constructing a user model based on partial experience with that user ”. Recent research highlights the importance of adaptable mHealth technologies in facilitating chronic disease self-management (McLean et al., 2011; Grua et al., 2020). An example are mobile applications developed for the management of diabetic hypoglycemia, which adapt the navigation system for hypoglycemic events (Pagiatakis et al., 2020). Under normal circumstances, the app displays a standard homepage for regular use (see Figure 1(a)). However, as depicted in Figure 1(b), access to other sections of the app will be restricted during a hypoglycemic event, and a prominent feature allows users to quickly contact emergency services. Jabeen et al. (2019) design an adaptable interface called PD-Helper to cater to the physical needs of Parkinson patients (see Figure 1(c)). The web page control section includes web elements that allow users to adjust font size, reload the page, open a new page, and navigate back to the main menu with a single tap on the screen.
Chronic Disease Self-Management: Chronic illnesses pose a considerable health issue on a worldwide scale, accounting for a significant number of yearly deaths (WHO, 2022). These diseases are responsible for the deaths of 41 million individuals annually, representing 74% of all global fatalities (WHO, 2023). The prevalence of chronic diseases is steadily increasing, driven by the reclassification of previously fatal diseases as chronic diseases and the ageing of the population (Beaglehole et al., 2008). Despite advancements in medical science, significant challenges persist in finding cures for chronic diseases such as diabetes and Alzheimer’s disease (Hird et al., 2016). Traditional treatment paradigms do not address the multifaceted nature of chronic diseases (WHO, 2022; Hird et al., 2016), as treatment cannot be based solely on biological parameters, which require active engagement and self-management by patients (Deiss et al., 2006). Self-management entails actively engaging in self-care activities to enhance one’s behaviour and well-being (Deiss et al., 2006). Research has highlighted the efficacy of mHealth technology in supporting self-management of chronic diseases (Paré et al., 2007; Hamine et al., 2015). Although there are possible advantages of mHealth interventions for individuals with long-term health problems, studies indicate a trend where those who could benefit the most from these interventions tend to have lower utilisation rates (Han et al., 2010). A significant proportion of fatalities resulting from chronic diseases, accounting for 77%, occur in low and middle-income nations (WHO, 2023).
AUIs and Healthcare: Previous research has explored the application of AUIs in systems adapted for healthcare professionals (Eslami et al., 2018; Greenwood et al., 2003; Vogt and Meier, 2010). Eslami et al. (2018) conducted interviews and observations to investigate user preferences regarding data entry, language and vocabulary, information presentation, as well as providing help, warning, and feedback, with a primary focus on healthcare professionals. Similarly, Vogt and Meier (2010) examined AUI design issues that aim to simplify input and reduce the potential for errors, particularly in contexts such as smart hospitals. Greenwood et al. (2003) introduced a novel approach using reactive agents for AUIs in diabetes treatment decision support, customising data display according to clinician preferences. Despite the initial emphasis on healthcare professionals, there is an increasing recognition of the diverse user base of mHealth applications. Consequently, there has been an increasing number of literature describing various approaches, applications, and tools specifically tailored for patient-focused AUIs. Existing studies on AUI frameworks often focus on particular adaptive components or specific aspects of patient management (Shakshuki et al., 2015; Fröhlich et al., 2009; Yuan and Herbert, 2012). Shakshuki et al. (2015) proposed an AUI architecture for patient monitoring with a focus on health-related information adaptation. Fröhlich et al. (2009) explored the application of LoCa, primarily to monitor physiological data and the activity status of patients within a digital home environment, facilitating context-aware adaptation of workflows. Yuan and Herbert (2012) designed a fuzzy-logic-based context model for personalised healthcare services in chronic illness, prioritising health problem prediction and preventive measures based on user data rather than UI adaptation to individual user needs and context.
Despite the initial emphasis on healthcare professionals, there is an increasing recognition of the diverse user base of mHealth applications. AUIs have been deployed across a spectrum of mHealth applications, ranging from stroke rehabilitation (Burke et al., 2009), diabetes (Pagiatakis et al., 2020), cardiac disease (Mohan et al., 2008), dementia (Awada et al., 2018; Gullà et al., 2019) and Parkinson’s disease (Jabeen et al., 2019). These applications exhibit varying degrees of adaptability, ranging from the adjustment of difficulty levels of exercise activity (Burke et al., 2009) to the customisation of health-related information (Awada et al., 2018; Gullà et al., 2019), navigation adaptations (Pagiatakis et al., 2020; Awada et al., 2018), multimodal interfaces (Pagiatakis et al., 2020), information architecture (Mohan et al., 2008), and graphic design (Gullà et al., 2019; Mohan et al., 2008; Pagiatakis et al., 2020; Jabeen et al., 2019). However, most of these studies lack detailed explanations of their AUI development process, especially concerning the initial stages involving the collection of diverse end user requirements. Furthermore, the typical evaluation approach for AUIs focuses on overall application effectiveness without proper comparisons to non-adaptive UIs, which complicates drawing specific conclusions about the impact of AUIs. Although research in other domains suggests that AUIs can improve user performance and satisfaction compared to non-adaptive baselines (Gajos et al., 2006; Paymans et al., 2004; Tsandilas and Schraefel, 2005; Li and Gajos, 2014), disruptive adaptations, which alter users’ accustomed interaction patterns or break conventions, can result in frustration or dissatisfaction (Findlater and McGrenere, 2004; Rivera, 2005). Despite these insights, our understanding of how individuals with chronic diseases use AUIs and how to design mHealth technologies integrating AUIs to maximise benefits while minimising costs for this population remains limited.
AUIs and Chronic Disease Self-Management: To enhance the successful deployment of mHealth applications, particularly for patients managing chronic conditions, it is imperative to tailor these applications to meet the needs of diverse user groups. There are well-recognised challenges to doing this, including: Firstly, chronic diseases have high heterogeneity, impacting patients differently in terms of triggers, symptoms, and severity (Harvey et al., 2012; Deiss et al., 2006; Audulv, 2013). Therefore, patients may have various self-management needs. Secondly, the design of UI must accommodate the changing phases of chronic diseases over time (di San Marco et al., 2019; Audulv, 2013). Many chronic conditions either worsen or improve with time, necessitating corresponding adjustments to self-management plans (Lorig and Holman, 2003). Furthermore, chronic diseases often co-morbid with other medical or psychological disorders, resulting in a broader spectrum of user characteristics and functionality requirements (di San Marco et al., 2019; Islam et al., 2014). For example, diabetes can lead to various complications such as vision loss, amputation, neuropathy, end-stage renal disease, cardiovascular disease, infections, and cognitive impairment (Vijan, 2015). Thirdly, the vast majority of chronic diseases are long-lasting and generally lifelong (WHO, 2022; Harvey et al., 2012), typically manageable but not curable. Therefore, mHealth technologies must sustain user engagement and motivation over the long term. In addition to the diverse nature of chronic diseases, patients also have a wide range of backgrounds, expertise, and demographic, psychological, and cognitive characteristics (Vasilyeva et al., 2005). While adapting interfaces can improve user acceptance and motivation (Duplaga, 2015; Han et al., 2010), existing AUI solutions often lack full adaptability, relying on predefined rules and overlooking unique characteristics of users (Grua et al., 2020; McLean et al., 2011). We wanted to explore the perspectives of users on AUIs in chronic disease management, seeking to establish guidelines for designing more effective AUIs that meet the varied needs of individuals with chronic diseases.
3. Methodology
We developed an AUI prototype of the mHealth application and gathered input from various users with chronic disease through a combination of focus group sessions, interviews, and surveys. Through the analysis of the data collected, we derived guidelines for effective AUI design for mHealth app targeting chronic disease. Subsequently, we refined these guidelines based on the feedback of the participants, as shown in Figure 2.
3.1. Adaptive User Interface prototype
Building upon insights from an earlier Systematic Literature Review (SLR) (Wang et al., 2023b), our research enhances the adaptation categories and integrates them into a single prototype that includes three primary types of adaptation: presentation adaptation, content adaptation, and behaviour adaptation. Details of the prototype can be found in (Wang et al., 2023a).
Presentation adaptation is a crucial aspect of UI design, aiming to enhance user experience by modifying the UI appearance. This adaptation involves adjusting elements such as colour, positioning, and font size to better suit the user’s preferences and needs. It is divided into two main categories: 1) graphic design: manipulates visual aspects, like theme, and layout; 2) information architecture: pertains to the structural organisation of information within the system.
Content adaptation is the process of adapting interface content to better suit user needs and preferences, encompassing various elements such as text, semantic elements, images and explanatory text. We propose two primary categories: 1) content complexity: simplifies content for better comprehension, which comprises three subcategories. The easy-to-understand language employs clear and concise language to convey information in a straightforward manner (see Figure 3(a)). Minimalist design streamlines the presentation of content by minimising visual clutter and extraneous details. Text-to-image conversion converts complex textual information into easily understandable visual representations, such as diagrams or illustrations, to aid comprehension, particularly for users who may struggle with textual content (See Figure 3(b)). 2) interface elements rearrangement: includes two main subcategories: position adjustments and selective element concealment. Position adjustments optimise the placement of interface elements to ensure intuitive navigation for users. The concealment of selective elements strategically hides or reveals interface elements based on user context or preferences, thus simplifying the interface and focusing on key content areas (see Figure 3(c)).
Behaviour adaptation involves adjusting navigation methods, enabling/disabling interface elements, and modifying interaction modes within an application. This complex form of adaptation often spans multiple steps and can encompass content and presentation adjustments. We group it into five subcategories: 1) navigation adaptation: alters user navigation permissions or suppressing specific modules; 2) add-on functions: incorporates new features to enhance user assistance and application usability; 3) pervasive strategies, adapts motivation techniques for effective behavioural changes tailored to user types/status; 4) multimodal interaction: adjusts interface modalities based on varied usage contexts; and 5) difficulty level adjustment: modifies game/exercise difficulty to balance engagement and challenge based on user motivation/performance. The AUI prototype was evaluated by three coauthors and tested on two real users with experience with mHealth apps. For each adaptation, we provided transitions, sample rationales, and additional instructions to facilitate user navigation and interaction with various adaptations.
3.2. User study design
Our user study consists of two parallel investigations, as illustrated in Figure 2. We conducted a qualitative investigation via focus groups and interviews to examine how individuals experience AUIs in the context of chronic diseases using the AUI prototype. At the same time, we conducted a quantitative survey to collect user preferences regarding different dimensions of adaptations.
3.2.1. Focus group and interview studies
We designed a focus group and interview protocol consisting of three sections. The first section collected detailed demographic information about participants, including their chronic disease and their usage pattern of mHealth apps through a Qualtrics survey. The second section collected the participant’s views on the different adaptations we present in the AUI prototype. The participants initially reviewed a brief adaptation video accompanied by audio explanations, providing detailed introductions to each adaptation. This visual aid was followed by hands-on interaction with the prototype. Instructions were provided in the accompanying slides, encouraging participants to actively engage with the prototype. In the event of any difficulties, participants could refer to the instructions for assistance, ensuring a seamless user experience. The researcher remained readily available to offer support, allowing participants to focus solely on the example adaptations. Given the interpersonal dynamics inherent in the interview and focus group settings, participants may tend to offer responses that they perceive as socially desirable (Harris and Brown, 2019). This potential bias is further compounded by the qualitative nature of the data analysis process, which is based on inductive coding and is susceptible to the subjectivity of the researcher (Hoda, 2021). We adopted between-method triangulation to address inherent flaws in each method. This involved the triangulation of data using a combination of quantitative and qualitative techniques (Fusch et al., 2018; Denzin, 2017; Oppenheim, 2000). In the third section, participants in the interview and focus group study were required to complete a post-session survey to express their preferences for various adaptations explored with the AUI prototype during the session. This post-session survey consists of 7 questions. Further details of the post-session survey are provided in Appendix A. As a token of appreciation, all participants were offered a AU$30 virtual gift voucher.
Both the focus group and the interview sessions were subjected to one round of pilot test that involved varying numbers of participants. We employed a theoretical sampling method (Hoda, 2021) for participant recruitment. This sampling approach involves selecting new participants based on specific criteria as data collection and analysis progress and yield categories and concepts, allowing the deepening or expansion of existing categories and concepts. Qualitative data collection and analysis followed an iterative process spanned three different data collection iterations (see Table 1). Upon obtaining ethics approval from the university ethics board, 111Monash Human Subject Ethics Committee approval number: 36568 we started our focus group study by reaching out to potential participants through a mailing list associated with a previous chronic disease research programme. In this iteration, our main focus was on recruiting older individuals (Iteration 1). Subsequently, using our personal connections and networks, we expanded our participant pool to include several young adults with chronic diseases (Iteration 2).222The first author, who is afflicted by a chronic condition, engaged with both online and offline support groups for individuals in similar circumstances. To broaden the inclusion of participants, we also used the snowball method (Parker et al., 2019) during Iteration 2. To improve the reliability and validity of the data through multiple data collection methods (methodological triangulation) (Denzin, 2017), we employed semi-structured interviews alongside focus group studies in Iteration 3. The group dynamics inherent in focus group research provides a unique perspective for observing collective attitudes and shared perceptions (Acocella, 2012). At the same time, interviews allow for personalised exploration of individual experiences and perspectives, fostering a detailed understanding of nuanced perspectives (Wilson, 2014). Interview participants were recruited through a multifaceted approach. Using local community channels, recruitment efforts were expanded to include prominent healthcare facilities such as the Baker Institute Diabetes Clinic and Alfred Hospital (Iteration 3).333https://kidney.org.au, https://strokefoundation.org.au, https://www.dementia.org.au, https://www.alfredhealth.org.au, https://baker.edu.au/health-hub/clinics/diabetes-clinics
3.2.2. Survey
Our post-session survey aims to gather additional data about user perspectives regarding our prototype and to triangulate them with the findings from the focus groups and interviews. In parallel, we conducted an anonymous online survey using Google Forms aimed at individuals with chronic diseases to gather feedback on their preferences regarding various aspects of collecting user’s preference for different aspects of adaptation. These two surveys differ in their question designs and serve for different purposes. The survey consisted of four main sections, each focused on specific topics to ensure participants’ ease of understanding: demographic information, health status, mHealth application use patterns, and AUI preferences. To ensure the participants’ understanding of AUIs, a brief explanation with two examples of AUIs in the health domain was provided at the beginning of the AUIs section. This section focuses on participants’ preferences regarding various adaptations, different types of data, data collection methods, and the desired level of participation in the adaptation process. The survey questions were informed by a previous SLR (Wang et al., 2023b), and the survey went through expert reviews and pilot tests with five participants. The survey data collection occurred from January to July 2023 through various recruitment channels, with a total of 90 respondents (Figure 2). The survey data collection went through three phases: 1) email ads and social media (58 responses), 2) physical posters distributed at Baker Institute Diabetes Clinic and Alfred Hospital (13 responses), and 3) advertisements through Dementia Australia, Stroke Foundation, and Kidney Australia (19 responses). Participants who completed the survey were eligible for a AU$20 virtual gift card draw. You can find the survey instruments in the Appendix B.
3.3. Guidelines Development
Our guideline development process follows and adapts the framework proposed by Hermawati and Lawson (2015) and Quiñones et al. (2018). We use data collected during survey, interview, and focus group sessions, as well as the existing literature on guidelines, design considerations, and criteria, to establish the guidelines. By synthesising this qualitative and quantitative information, we were able to identify common categories, patterns, and user needs that served as the basis for our guideline development. In addition, our guideline specification involved a process of refinement and validation, where we refined the guidelines based on the feedback of the participants. After the refinement process, we generated nine guidelines that serve as a practical and effective resource for designing adaptive mHealth applications targeting chronic diseases.
3.4. Data analysis
3.4.1. Qualitative data analysis
We used the data analysis procedures of Socio-Technical Grounded Theory (STGT) (Hoda, 2021) to analyse the data from the focus group and interview study. This decision was primarily driven by two key factors. First, the focus of our study closely aligns with the principles of the socio-technical research framework that STGT is built upon. Our investigation revolves around AUIs in applications related to chronic diseases, a socio-technical phenomenon that encompasses both human and social aspects (such as chronic disease, ageing, and computer literacy), as well as technical aspects. Second, STGT allows for selective application by integrating its core data analysis procedures of open coding (with hashtags) and constant comparison, along with memoing, within a mixed-method research study. Traditional grounded theory methods such as Glaserian (Glaser, 1992) and Strauss-Corbinian (Strauss and Corbin, 1994) are developed as standalone methodologies for theory development. By leveraging STGT for data analysis, we adopted a holistic approach to data analysis, recognising the complex and interconnected nature of the socio-technical phenomena under investigation. We obtained consent from the participants to transcribe the audio recordings, and subsequently stored and analysed the data using NVivo.
The data collection and analysis process followed an iterative and interleaved approach, as depicted in Figure 4. Saturation, which indicates the point at which no new categories or concepts properties emerged, was achieved by Iteration 3. The qualitative data was analysed by the first author and was subsequently shared with the remaining authors to encourage collaborative discussion at each stage of the process. This collaborative effort helped refine the analysis and determine the optimal presentation of the findings in all iterations of data collection. In Section 4, we present the key concepts and categories derived from the STGT analysis. An example of process of applying STGT for data analysis (Hoda, 2021) is provided below:
Open Coding and Constant Comparison: We analysed the audio transcripts and extracted various codes from the raw data. We provide some examples below.
Raw Quote: ”If you want to set it, then you don’t want to change it. Likewise, you don’t want to have to go through all those settings again. I think that would be complicated, and I am not sure I could handle it on my own. ” (FP 4)
Code 1: mental workload for adaptationRaw Quote: “I am not going to log my daily blood sugar levels in the app because I am too lazy to do it sometimes. Will the adaptation hide this important function because I don’t use it a lot?” (FP 6)
Code 2: user preferences contradict the app’s intended usageConcept: The two code examples given above suggest a concept: ”The user may not be the right person to handle the adaptation”
Category: Drawing insights from the memos generated during the coding process and the codes and concepts identified, we derived various categories. The given example illustrates the category concerning who should take charge of the adaptation process.
Theoretical sampling: We started our initial convenience sampling with the primarily elderly in three focus group studies Iteration 1. After we conceptualised concepts from the focus group data, we applied theoretical sampling to obtain more data that helped us to explicate the emerging concepts. For example, most of the focus group participants, who are older, express difficulty engaging in the adaptation process and prefer it to occur without their direct involvement, as the multistep nature of the process can be confusing for them. Conversely, a participant of a much younger age has a contrasting view, showing interest in adapting the features or functions. As a result, the category ”The user may not be the right person to handle the adaptation” . needs to be refined, prompting us to recruit more participants from a younger age group in Iteration 2 (see Figure 4).
Memoing played a crucial role in our approach, allowing us to explore emerging concepts and potential relationships between them, as outlined by Hoda (2021). These memos served as invaluable tools for capturing key insights and reflections gleaned from our open-coding efforts, which are further detailed in Section 4.5.
3.4.2. Quantitative data analysis
We exclusively gathered quantitative data through our survey, and subsequently performed statistical analysis using Microsoft Excel. Descriptive statistics provide a comprehensive overview of the characteristics and preferences of the survey respondents, allowing us to understand the distribution and trends within the dataset. Chi-square tests were employed to determine if there was a significant association between the preferences of users with respect to various aspects of adaptations and their demographic characteristics, such as age, gender, nationality, education, and chronic diseases. To ensure that the data distribution meets the prerequisites for the Chi-square independence test, related variables were grouped into categorical variables beforehand. Age was categorised into two groups: 18-45 and 45-74. The level of education was classified into three groups: Less than Bachelor’s degree, Bachelor’s degree, and Postgraduate (Master’s and Doctoral degree). Chronic disease conditions were categorised as detailed in Section 5.1. If a significant association is found, binary logistic regression or multinomial logistic regression will be subsequently employed to model the relationship between these variables. Understanding how these demographic factors influence user preferences provides valuable insights into tailoring AUIs to accommodate diverse user needs and preferences. The common significance level of is chosen for statistical analysis.
4. Findings from interviews and focus groups
We present our findings on user’s perspectives towards the AUI prototype. We identified four overarching challenges that participants faced while interacting with our AUI mHealth prototype (outlined in Table 2) . We also found a set of strategies applied or purposed by participants to have a better user experience of the adaptations (outlined in Table 3). These strategies were founded to be moderating by different characteristics of users.
| # | Age | Gender | Ed* | Nationality | Chronic diseases | mHealth app usage pattern |
| Iteration 1: Each focus group study took 60-70 minutes. | ||||||
| FP1 | 55-64 | Female | M | Australian | Type 2 diabetes | Less than once a week/1-10 minutes |
| FP2 | 55-64 | Female | M | Australian | Type 2 diabetes | Less than once a week/20-30 minutes |
| FP3 | 65-74 | Female | M | Australian | Type 2 diabetes | Once a week/10-20 minutes |
| FP4 | 65-74 | Male | P | Australian | Type 2 diabetes | 2 or more times per day/1-10 minutes |
| FP5 | 25-34 | Female | M | Chinese | Asthma, Allergic rhinitis & Chronic gastritis | Less than once a week/1-10 minutes |
| FP6 | 65-74 | Female | P | Tamil | Type 2 diabetes | Less than once a week/1-10 minutes |
| FP7 | 55-64 | Female | T | Australian | Type 2 diabetes | Less than once a week/1-10 minutes |
| FP8 | 18-24 | Female | B | Malaysia | Depression | Less than once a week/1-10 minutes |
| FP9 | 65-74 | Male | B | Australian | Type 2 diabetes, Heart disease, Epilepsy&Asthma | 2 or more times per day/1-10 minutes |
| Iteration 2: Each focus group study took 60-70 minutes. | ||||||
| FP10 | 25-34 | Male | M | Chinese | Type 2 diabetes | 2 or more times per day/1-10 minutes |
| FP11 | 25-34 | Female | B | Filipino | Over weight, Asthma & Anxiety | 2 or more times per day/1-10 minutes |
| FP12 | 25-34 | Male | M | Chinese | Rhinitis & Anxiety | 2 or more times per day/1-10 minutes |
| FP13 | 35-44 | Male | B | India | Overweight & Type 2 diabetes | less than once a week/1-10 minutes |
| FP14 | 35-44 | Male | B | India | Hypertension | less than once a week/1-10 minutes |
| FP15 | 25-34 | Female | M | Chinese | Mental health conditions | 2 or more times per day/more than 30 minutes |
| Iteration 3: Each interview study took 30-60 minutes (aimed at saturation). | ||||||
| IP1 | 55-64 | Female | A | Australian | Type 2 diabetes | 2 or more times per day/1-10 minutes |
| IP2 | 35-44 | Male | B | British | Mental health conditions | Once a week/1-10 minutes |
| IP3 | 65-74 | Female | A | Australian | Type 2 diabetes | 2 or more times per day/1-10 minutes |
| IP4 | 45-54 | Female | M | Australian | Hypertension& Type 2 diabetes | 2 or more times per day/1-10 minutes |
| IP5 | 45-54 | Female | M | Chinese | Anxiety, Insomnia & Mental health conditions | 2 or more times per day/1-10 minutes |
| IP6 | 18-24 | Female | B | American | Anxiety & Mental health conditions | 2 or more times per day/1-10 minutes |
| IP7 | 55-64 | Female | B | Australian | Major chronic depressive disorder & PTSD | 2 or more times per day/1-10 minutes |
| Ed*: Education ( M=Master, B=Bachelor, A=Associate, T=Technical training ) | ||||||
4.1. Focus group and interview participants
Table 1 provides a breakdown of the participant characteristics, showcasing a variety of nationalities such as Chinese, Indian, British, Filipino, Malaysian, American, and Tamil, with a notable majority being Australian (n=10). The sample comprises a greater proportion of females (n=15) than males (n=7), spanning across various age groups (18-34: 7, 35-54: 6, 55-74: 10), as well as a wide spectrum of chronic diseases, such as type 2 diabetes (n=12), rhinitis (n=2), chronic gastritis (n=1), epilepsy (n=1), hypertension (n=2), asthma (n=3), chronic depressive disorder (n=1) and different mental health conditions (n=7). All participants have current or previous experience with mHealth applications. Their experience is assessed based on the frequency and duration of their usage (Robbins et al., 2017). Participants were asked to indicate how frequently they use health apps and to specify the duration of each session. The result indicates varying usage patterns among participants. 8 users reported using the app less than once a week, while 12 mentioned using it two or more times a day. Most users (n=19) reported a session duration of 1-10 minutes. Most of the participants (n=17) hold a university degree. Of these, almost half have obtained a bachelor’s degree (n=8), and the remaining half have earned a master’s degree (n=9). Participants are numbered according to their study involvement: FP1-FP15 for focus group participants and IP1-IP7 for interview participants.
4.2. Challenges participants face in using AUIs
Drawing on STGT data analysis and the existing literature (Wang et al., 2023b; Abrahão et al., 2021), we grouped the identified challenges into four distinct categories underpinned by the key concepts (see Table 2). These categories are as follows: (1) What to adapt: which revolves around identifying the specific elements or aspects of the UI that require adaptation. It involves understanding the potential challenges that may arise when different types of adaptations are implemented; (2) Who should initiate adaptations: This category addresses the roles and responsibilities associated with initiating interface adaptations; (3) How to adapt: This category focuses on the mechanisms and strategies utilised to effectively implement interface adaptations; (4) When to adapt: This category pertains to determining the optimal timing and conditions under which adaptations should take place.
| Categories | Concepts | Participants | # | Moderating effect(NA:Not Applicable) |
| C1. What to adapt (10 FP, 7 IP) | Predictability of the user interface | FP 2, IP 4,5,7 | 4/22 | NA |
| Comprehensibility of adaptation | FP 2,4,7,9,12,15,16, IP 3 | 8/22 | Technological proficiency | |
| Obtrusiveness of generated adaptations | FP 4,5,6,7,9, IP 1,3,4,6,7 | 10/22 | mHealth usage experience | |
| Mismatch of generated adaptations | FP 4, IP 2,3,6 | 4/22 | Self-sufficiency and independence | |
| Aversion to negative measurements | FP 8,9, IP 1-5, | 7/22 | Health conditions&Coping mechanisms | |
| C2. Who should initiate adaptation (12 FP, 6 IP) | System is unable to initate the adaptation | FP 4-6,9,11,12,15, IP 3,6 | 9/22 | NA |
| User might not be the right person to handle the adaptation | FP 1-4,6,7-9,11,12, IP 1-4,6,7 | 16/22 | Cognitive load | |
| C3. How to adapt (4 FP, 3 IP) | Infringement of privacy | FP 11,13, IP 2 | 3/22 | NA |
| Decision fatigue and frustration | FP 7,9, IP 2-4 | 5/22 | NA | |
| C4. When to adapt (6 FP, 6 IP) | Frequency of adaptation | FP 4,6,8,9, IP4,7 | 6/22 | Characteristic of features |
| Timing of adaptation | FP 4,6,7,8,10, IP 1,3,4,6 | 9/22 | Health conditions&mHealth usage experience | |
| Usage scenario | FP 8,9,10, IP 1,3,4,5,7 | 8/22 | NA |
4.2.1. C1. What to adapt
The importance of UI predictability emerged as a significant concern, consistent with prior research findings (Mezhoudi et al., 2015; Jameson, 2007). The participants emphasised the importance of spatial stability within the UI (IP 4,5,7, FP 2). Participants rely on familiar patterns and locations within the UI to locate information and perform tasks efficiently. IP7 articulated, ”system makes lots of changes would confuse me because I get used to finding things in the same place. When it changes, it is hard to see what will come next”. Studies have shown that even minor changes in interface layout or functionality can disrupt users’ mental models and increase the likelihood of errors (Hammond, 1987). Therefore, ensuring stability and predictability in UI design is essential for supporting users’ cognitive processes.
Some participants faced difficulties in understanding the concepts of UI adaptation and the resulting adaptations, despite the assistance and explanations provided by the interviewer or facilitator (FP 2,4,7,9,12,15,16, IP 3). This lack of comprehensibility could potentially hinder the effectiveness of adaptations, leading to a decrease in user performance and satisfaction, and ultimately causing users to consider switching to alternative applications (FP 2). Within the context of chronic diseases, the intricacies of adaptation may be further exacerbated (Yuan and Herbert, 2012). For example, in our AUI prototype, an avatar displayed a concerned expression in response to unfavourable measurements. However, older participants (over 65 years) (FP 4, IP 3) suggested that this simple visual cue may not alleviate their concerns and may even increase their confusion. On the contrary, younger participants (aged less than 44 years) appreciated such avatars, viewing them as a means to ”make all serious things more fun” (IP 7, FP 5, 8, 10, 13). This divergence is consistent with previous research indicating that gamified design is more appealing to younger audiences due to their technological proficiency and familiarity with gaming (Betts et al., 2019). The complexity inherent in the management of chronic diseases requires adaptation strategies that are easily understandable for users of different age groups and backgrounds. Older adults may require straightforward and explicit communication of adaptations. On the contrary, younger users may appreciate more interactive and engaging elements.
Some participants expressed concerns about specific adaptations, particularly those that involve comfort words and gamified design elements to convey health-related information, citing them as obtrusive and potentially diverting attention from primary tasks (IP 1,3,4,6,7, FP 4,5,6,7,9). For example, instead of simply stating the health measurement result in a clinical way, the app could say something like: ”I’m sorry to hear about your recent health measurement result. Remember, setbacks are a normal part of the journey. Let us work together to find ways to improve your health”. Additionally, it is important to consider the broader challenges associated with gamification in mHealth applications, including issues related to gamification appropriateness and socio-cultural adaptability (Vajawat et al., 2021). Gamification affects patients differently, and mismatched adaptations may have a negative influence on patients (Codish and Ravid, 2014). Participants with more experience using mHealth applications tended to view certain distracting adaptations less negatively (FP 5, IP 6). This suggests that familiarity with the mHealth tools may influence how users perceive and interact with various design elements (McDougall et al., 2000).
Mismatched adaptations were also mentioned a lot by the participants (FP 4, IP 2,3,6). They worried about being trapped in an adaptation process triggered by ”wrong measurements or incorrect input data” (FP 4, IP 6). IP2 shared similar feelings, stating that he ”felt overwhelmed due to the uncertainty of what the system would understand and what it would not”. Such instances of expectation mismatches often result in a decrease in trust in the adaptation generated (Glass et al., 2008). Integration of sensor networks and data in mHealth has the potential to improve healthcare services for patients (Albahri et al., 2018). However, instances of ”sensor failure” (IP 3) or ”inaccurate input data” (IP 6) from patients can raise concerns about the generation of unwanted adaptations. Furthermore, some younger participants (aged under 34) acknowledged the significance of certain adaptations, such as adjustments in font size or simpler UI designs (FP 8,11,15, IP 6). However, they often perceived these adaptations as beneficial not for themselves but for individuals with varying disabilities and levels of independence. Their hesitation to fully embrace these adaptations may be attributed to their strong sense of self-sufficiency and independence (Trajkova and Martin-Hammond, 2020), which leads them to overlook the immediate advantages of such adaptations. These perspectives highlight the importance of focusing on long-term benefits or positive impact on individuals, broadening the acceptance of adaptations.
Participants in our study exhibited mixed feelings about the measurement results, particularly in the context of recording health data such as blood glucose levels, stress levels, and blood pressure. Some participants expressed a strong aversion towards bad measurement results and preferred not to be constantly reminded of negative outcomes (IP 1-5, FP 8,9). For example, IP1 shared her tendency to avoid doctors and attempted to deny the measurement when faced with unfavourable health data. Participants suggested that presenting the data in a ”noncritical” (IP 4) and ”positive reinforcement” (IP 3) manner would be beneficial. These struggles highlight the influence of user symptoms and health conditions on their interactions with technologies (Kornfield et al., 2020). However, other participants, particularly those with mental health problems, found value in reminders and clear presentations of such data as important signals for their health management (IP 5, FP 5,8,10-13,15). For example, IP5, who deals with mental health issues, expressed that:
The different attitudes towards measurement results observed among participants suggest different coping mechanisms employed to manage stressful situations (Csikszentmihalyi et al., 2014). Some individuals may adopt avoidance and denial as coping strategies, others may employ more positive coping strategies to reduce stress and help the mind and body return to their original state of functioning (Folkman and Lazarus, 1980).
4.2.2. C2. Who should initiate adaptation
In typical app usage scenarios, users already face the challenge of continuously recording various types of data, such as heart rate, blood pressure, blood sugar levels, and body composition (Philip et al., 2022). The introduction of AUIs in these applications can further complicate matters, primarily due to the crucial role of data in driving UI adaptation (Alvarez-Cortes et al., 2007). Many participants in our study already struggle to maintain complete and accurate daily data records, resulting in occurrences of incomplete data that impede the system’s ability to initate the adaptation process (FP 4-6,9,11,12,15, IP 3,6). For example, FP9 revealed his frequent lapses in recording blood tests, while FP5 acknowledged that he forgot to record her daily mood despite setting up multiple reminders. Users may struggle to enter their health data for various reasons, including forgetfulness, lack of motivation, or difficulty integrating the use of the app into their daily routines (Stawarz et al., 2014). This inconsistency in data recording is a common issue in mHealth applications related to chronic diseases (Short et al., 2018). This problem is often attributed to the low levels of user adherence and the high rates of attrition commonly observed in mHealth interventions (Kelders et al., 2012). Furthermore, IP3 admitted that sometimes ”fakes some data” to complete records artificially, which can ”make the progress and reports look really good”. These participants emphasised that having various adaptations becomes pointless if the underlying data is unreliable. As a result, the reliability and completeness of health records can be compromised, limiting the adaptation of these functions in monitoring and managing chronic disease functions.
Based on the discussion of data collection, participants expressed concerns about whether user is the right person to handle the adaptation process. Firstly, older participants (over 45 years) mentioned the added mental workload of overseeing adaptation processes along with their existing responsibilities related to managing their chronic conditions (IP 1,4,7, FP 4,6,7,9). This additional cognitive load could potentially overwhelm them, hindering their ability to effectively self-manage their health. IP1 noted that:
Added to this, participants also found it challenging to learn and navigate the adaptation process, which could counteract the time and effort savings benefits that automation of adaptations aims to offer (Gajos et al., 2006; Harman et al., 2014). Some participants expressed reservations about complete automation, fearing that it could lead to a lack of awareness about system changes (IP 3,6, FP 1,2,3). This suggests that mHealth apps could pose risks to user safety when their design is not well controlled (Akbar et al., 2020).
Secondly, considering the intended purpose of the app, situations can arise where a user’s preferences conflict with the intended use of the app (FP 4,6,7-9,11,12, IP 2). For example, in the context of rearranging elements on the UI, FP4 expressed concerns that:
FP11 and FP12 suggested that some users might manipulate the system to obtain the adaptations they want, deviating from the original design and purpose of an app. For instance, a user who always forgets to record or report lower activity levels than the actual, hoping to receive suggestions for lighter workouts or fewer reminders to be active. In these cases, users manipulate the system to obtain adaptations that deviate from promoting genuine physical activity and supporting users’ health goals.
4.2.3. C3. How to adapt
mHealth devices are capable of monitoring various aspects, including activities, sleep patterns, emotional states, vital signs such as blood pressure, and mental health conditions (Prasad et al., 2014). Several younger participants (under 44 years of age) in our study raised concerns about potential privacy infringements related to adaptation (IP 2, FP 11,13). IP2 emphasised that he ”probably would not allow any audio-based input for every app” because he was worried that the app is listening all the time. This observation implies that smartphone users may feel violated when discovering that applications access their data without their knowledge (Li et al., 2020), potentially leading individuals with greater privacy concerns to avoid using such apps (Lohr, 2015). On the contrary, other participants expressed confidence in the reputation of the application (FP 1,2,3,4,6) or the organisation (FP 14,15, IP 1,3,5,7) that developed the application, suggesting that they are not overly concerned about privacy issues. This sentiment is deemed reassuring given the prevalent privacy issues surrounding mHealth apps (Haggag et al., 2022). Furthermore, participants stressed the importance of implementing stringent data protection regulations and accountability mechanisms to prevent unauthorised use or disclosure of user data (FP 14,15, IP 2).
Excessive control over adaptation can result in decision fatigue and frustration, especially concerning individuals with chronic diseases already facing making numerous decisions related to their treatment and lifestyle adjustments (Scherer et al., 2021). Requesting participants to decide on adaptations might make them feel more drained as the day progresses (IP 2,3,4, FP 7,9), as this forces them to make more decisions (Sara, 2021). Stress associated with making frequent adaptation decisions can exacerbate existing health conditions and negatively impact overall well-being (Scherer et al., 2021). Participants indicated a preference for a ”system-driven approach” to adaptation decisions (IP 2).
4.2.4. C4. When to adapt
While the participants appreciated the ability to adapt the app, they found it challenging to engage with frequent adaptations (FP 4,6,8,9, IP 4,7), especially if they needed to perform these operations daily. FP6 expressed concerns about the time and effort required to manage constant changes, stating that it became overwhelming to keep up with the app’s adaptations alongside other daily responsibilities. Moreover, certain participants wished the UI to remain unchanged after the initial setup, particularly for less frequently used features (FP 4,6). They acknowledged the necessity of dynamic adjustments for frequently accessed features.
The timing of adaptation is a crucial aspect of UI adaptation (Bunt et al., 2004; Abrahão et al., 2021), and participants in our study expressed varying preferences in this regard. Some participants who were knowledgeable about their health conditions and app usage preferred setting the adaptation after logging in (IP 6, FP 8,10). This aligns with the notion that upfront customisation is efficient when users have a clear understanding of their needs (Bunt et al., 2004). On the contrary, the majority of participants preferred adaptations to occur at regular intervals but should wait until they had familiarised themselves with the application (FP 4,6,7, IP 1,3,4). This suggests that some users prefer a gradual introduction of adaptations to allow them to acclimate to the app’s interface and functionality before making any significant changes. FP7 expanded on this point, emphasising that immediate adaptation upon initial login could be overwhelming, particularly for those who had just been diagnosed with chronic diseases. Understanding user decision-making strategies is crucial, as they are influenced by their level of expertise in a given domain (Bettman, 1971). Expert users, who possess a deeper understanding of the app’s functionalities, can leverage their domain knowledge to exercise more precise control over the system. This highlights the importance of providing support to novice users while offering flexible customisation options for those with greater expertise. Such approaches can empower users of all levels to effectively navigate and leverage the adaptive capabilities of the app (Attfield et al., 2011).
Being attentive to various usage scenarios is essential to ensure that the application remains relevant and beneficial to a diverse user base (Stephan et al., 2017; Varshney, 2014). Understanding how users interact with the app in different contexts, such as when multitasking or when used by family members, allows developers to tailor the user experience to effectively accommodate these varied needs. For example, some participants noted that they could use the app together with their family members (FP 8,9,10, IP 1,3,4,5,7). In these cases, users have to ”look over the health data”, ”understand the features of the app”, and ”set up some reminders in the app” for their family members. This additional layer of complexity emphasises the multifaceted need for generated adaptations. IP7, for example, expressed the desire to understand why specific adaptations are made for their parents, which might not align with the parents’ priorities of using the app. Similarly, IP5, FP8, and FP10 mentioned the need to monitor their parents’ health while their parents are in a different country. This usage pattern in which the app is used by individuals other than the intended user, can potentially lead to unforeseen consequences (Lewis and Wyatt, 2014), such as misinterpreting data, incorrect health decisions, and feature misuse, especially when the primary user’s health needs differ from those of the secondary user (Aljedaani et al., 2021; Lewis and Wyatt, 2014). This interaction between users and potential users, including informal carers (e.g. children, spouses) and formal carers (e.g. doctors, nurses), is also highlighted in the study by Martin-Hammond et al. (2019).
4.3. Strategies to mitigate AUI challenges and relevant moderators
Participants discussed strategies relating to controllability, user support, and alignment need for designing adaptations during our user study session (summarised in Table 3). Participants exhibited varied priorities for different strategies, influenced by their preferred level of involvement in the system, their experience with mHealth technologies, and their own health conditions (Moderators).
| Categories | Concepts | Participants | #(%) | |
| Strategies | S1. Controllability and Autonomy (9 FP, 7 IP) | Dedicated adaptation panel to manage all adaptations | FP 7, IP 3,7 | 3/22 (14%) |
| Dedicating a specific part of the interface for adaptation | FP 4,6,9, IP 1,4,5 | 6/22 (27%) | ||
| Step-by-step adaptation process | FP 3,10,11,12, IP 1,3 | 6/22 (27%) | ||
| Configuring factors for different adaptations | FP 3,7,9, IP 2,3,4-7 | 9/22 (41%) | ||
| Accessible opt-in and opt-out process for the adaptation | FP 3,5,7,11,12,15, IP 1,3,4,6,7 | 11/22 (50%) | ||
| S2. User support for adaptation (6 FP, 6 IP) | Shortcut to different adaptations | FP 10,11,13, IP 2,3 | 5/22 (23%) | |
| Onboarding processes, educational materials & tutorials | FP 1,7,9, IP 1,5,7 | 6/22 (27%) | ||
| Provide user resource to personalise their experience | FP 7,9, IP 3,4,7 | 5/22 (23%) | ||
| Contextual explanations of adaptations | FP 7,8,9,13, IP 3,4 | 6/22 (27%) | ||
| S3. Alignment (6 FP, 10 IP) | Alignment of user’s chronic diseases | FP 1,2,3,4,6,9, IP 1,3 | 8/22 (36%) | |
| Alignment of user’s motivation and goal | FP 5,8,13, IP 3,4,7 | 6/22 (27%) | ||
| Alignment of existing technology and guideline | FP 3,4,6,7, IP 1,4,6 | 7/22 (32%) | ||
| Moderators | M1. User involvement | Active/passive user involvement | All participants | 22 (100%) |
| User-caregiver collaboration | FP 8,9,10, IP 1,3,4,5,7 | 8/22(36%) | ||
| M2. User experience with mHealth applications | All participants | 22 (100%) | ||
| M3. Health condition | All participants | 22 (100%) | ||
4.3.1. S1. Controllability and Autonomy
Participants in the study stressed the importance of having control and autonomy over the adaptation process. While autonomy and controllability are distinct concepts, they are closely related and can influence each other in the context of user interactions (Jameson, 2007). Some participants expect to have a dedicated adaptation panel that allows them to manage all adaptations (FP 7, IP 3,7). In this panel, they envision having comprehensive control over various aspects of adaptation, enabling them to fine-tune their user experience. For instance, IP3 explained:
In contrast to controlling adaptation in multiple places on the UI, some participants preferred to have a specific part of the system dedicated to adaptation which introduces only minimal distortion to the original interface (FP 4,6,9, IP 1,4,5). This preference resonates with the insights gained from previous studies (Gajos et al., 2006; Li and Gajos, 2014; Deuschel and Scully, 2016), possibly arising from the inherent desire to maintain high levels of spatial stability in the UI (Deuschel, 2018). By centralising adaptation controls in a specific part of the system, users can efficiently manage and customise adaptations without compromising the integrity of the overall interface (Deuschel, 2018). IP1 explained that she would feel more confident exploring the system’s adaptation features if there was a designated adaptation section within the app, as it would prevent ”accidentally making changes to different parts of the app”. This concern about unintentional changes is not uncommon among some elderly users (mentioned by FP 4,6,9), who may feel anxious about unintentionally altering the adaptation settings, often requiring assistance to address such issues. The confidence older individuals have in their technological proficiency significantly influences their adoption of mHealth (Cajita et al., 2018; Bentley et al., 2014).
One challenge previously highlighted is that users may not initially know which adaptation suits them best. Participants identified a potential solution by providing a step-by-step adaptation process that gradually introduces more complex adaptations (IP 1,3, FP 3,10,11,12). This approach aligns with research indicating that people tend to choose tasks of intermediate difficulty, which can improve their sense of competence (Deci and Ryan, 2013). However, it is essential to adapt in an unobtrusive manner that remains easy for users to notice and follow, as some participants expressed concerns regarding the engagement of older individuals (aged over 65) with frequent adaptation (FP 4,6). These different kinds of controllability and autonomy are achieved on various levels of granularity which vary from broad UI presentation changes to more fine-grained adaptations that impact dialogues, navigational flows, or specific content elements (Abrahão et al., 2021; Cristea and Calvi, 2003; Findlater and McGrenere, 2010). Our study uncovered a spectrum of preferences for the level of adaptation granularity. Some favoured one-time adaptations that encompassed the entire application, while others wanted specific adaptations targeting particular interaction objects.
The participants emphasised the need for flexibility in configuring multiple factors for different adaptations (IP 2,3,4-7, FP 3,7,9). Offering ”user control over the automation level and scope of adaptations” (IP 3), ”capability to prioritise UI elements based on their frequency, significance or sequence of usage” (IP 2), and ”options to mark specific adaptations as favourites for increased use or for similar adaptation scenarios” (FP 9) are noted by participants. Providing users with an easy and accessible opt-in and opt-out process for adaptation is crucial to ensure user autonomy and positive user experience (IP 1,3,4,6,7, FP 3,5,7,11,12,15). Participants desire a system that allows them to easily enable adaptations with a simple click (FP 3,7,11, IP 4), while also valuing the ability to revert adaptations if needed (FP 3,5). These practices for enhancing controllability empower users to maintain control over their app experience, they are more likely to remain engaged with the application over time, leading to greater user satisfaction and improved overall outcomes (O’Brien and Toms, 2008).
4.3.2. S2. Providing user support for adaptation
It is essential to provide comprehensive user support and resources to help navigate and understand adaptations. We identified two phases for this support: prior to adaptation and during adaptation. In the phase, before adaptation, users should have access to clear and detailed information about available adaptations, their potential benefits, and how to enable or disable them. Participants expressed the need for the system to offer straightforward and easily accessible shortcuts for making these adaptations, avoiding navigating through complex menus (IP 2,3, FP 10,11,13). For example, IP2 shared his experience of not noticing the adaptation feature in a mental health app for a year because it is buried deep within the settings menu. Additionally, incorporating onboarding processes or interactive tutorials can effectively introduce users to the concept of adaptation, demonstrate its benefits, and facilitate a better understanding of the underlying principles and purposes of adaptations (FP 1,7,9, IP 1,5,7).
During the adaptation process, users may need assistance or guidance to ensure smooth implementation. The availability of ample resources and meaningful suggestions plays a significant role in the desirability of adaptations. Participants have shown their appreciation for the option of selecting from a catalogue of adaptations offered by the application (IP 3,4,7, FP 7,9). As IP7 puts it:
This approach ensures that users have access to relevant suggestions while still maintaining the flexibility to adapt the technology to their preferences. Furthermore, providing contextual explanations for adaptations within the application to help users understand the adaptation process and, interpret perceived information is essential for transparency in the adaptation process (Höök, 2000). Participants frequently expressed the need for a detailed justification regarding the relationship between user data and generated adaptations, as well as the reasons behind why the system believes that a particular adaptation is needed or desired (IP 3,4, FP 7,8,9,13). Some participants also anticipated features such as ”highlighting the adapted parts on the interface” (IP 3) or ”using animated transitions to visualise the adaptation process” (FP 8). These thoughts on adaptation have been explored in prior research (Findlater and McGrenere, 2007) and reflect the importance of transparency in the operation of mHealth applications (Glass et al., 2008; Martin-Hammond et al., 2019). For instance, FP13 emphasised the importance of being able to ”verify the actions and outcomes for adaptation” as a primary factor driving his usage of the system. However, it is important to acknowledge that not all participants share this perspective. IP2, for instance, believed that contextual reasoning for adaptations is unnecessary if users have no control over what is suggested. This finding resonates with the conclusions drawn from the study by Bunt et al. (2012), which revealed that users exhibited minimal interest in explanations when they lacked direct control over an agent’s reasoning process. Furthermore, IP7 and FP4 cautioned against overly complex reasoning, which they believed could ”confuse and distract users from their primary tasks”. This divergence in opinion highlights the need to tailor support and explanations to users’ interests and cognitive abilities (Glass et al., 2008).
4.3.3. S3. Alignment
Achieving alignment in chronic disease management involves standardising and simplifying medical terminology and self-management processes. Older participants (over 55 years old) expressed the need for the system to provide a standardised, non-academic terminology for drug names (FP 1,2,3,4,6,9, IP 1,3). This request arises from the challenge of remembering the names of medications, particularly when managing multiple chronic conditions with frequently changing prescriptions. For example, FP9, who deals with heart disease, type 2 diabetes, epilepsy, and asthma, takes several medications concurrently. He faced challenges with the inconsistency in the appearance and names of medicine. FP1, FP3, and FP7 expressed similar feelings, asking for a simple and standardised naming system to aid memory recall. Additionally, some participants expressed a desire to acquire knowledge of medical terminology to facilitate more informed conversations with healthcare providers (IP 3,7). This indicates the need for a balanced approach that simplifies terminology for better comprehension while also offering avenues for users to broaden their medical understanding. The introduction of AUIs might inadvertently result in users acquiring less domain-specific knowledge compared to nonadaptive systems (Lanier, 1995). Thus, it becomes crucial to assess users’ readiness to compromise on knowledge acquisition in favour of ease of use when formulating AUI designs. Such an approach ensures that users can navigate complex medical information effectively, while also encouraging continuous learning and engagement with the application. Furthermore, many participants (FP 1,4,6,7,9, IP 1,3) recommended the incorporation of relevant visuals related to the app’s topic to enhance user understanding and engagement. For example, IP1 proposed incorporating a doctor’s image to signify critical health issues, suggesting the potential effectiveness of visual cues in conveying important health information. However, it is essential to recognise that the introduction of anthropomorphic visuals, such as the doctor’s image, may inadvertently shift users’ perceptions of responsibility away from themselves and onto the system (Höök, 2000).
Recognising the role of perceived value in user motivation, it is essential to ensure that mHealth applications offer tangible benefits aligned with users’ needs. Even with adaptation features, if users do not perceive practical relevance, sustained engagement becomes unlikely (O’Brien and Toms, 2008; Short et al., 2018). Some participants noted that while certain adaptations were convenient, they did not significantly enhance their chronic disease management (FP 5,8,13, IP 3,4,7). A participant remarked that ”people motivated solely by adaptation might be challenging to retain as users” (IP3), underscoring the importance of providing substantial value in mHealth apps for effective chronic disease management (Trajkova and Martin-Hammond, 2020; Melenhorst et al., 2001). Incorporating adaptations without clear utility or alignment with user goals can lead to confusion, increased cognitive load, and less intuitive interfaces.
To increase user engagement and familiarity, it is crucial to integrate applications with technology that users are already familiar with to enable simple and natural interactions. Some participants suggested incorporating voice assistants such as Amazon Echo or Google Home (FP 3,4,6,7, IP 1,4,6) as a means of providing an intuitive user experience. Furthermore, the use of established icon sets and standardised symbols can help users quickly understand the intended message, as discussed by IP2,4, and FP1,9. Participants often compared the design of the app with the familiar aesthetics and conventions of contemporary smartphone interfaces, especially in cases where users were less accustomed to encountering UI adaptive features (McDougall et al., 2000). This comparison reflects users’ expectations based on their prior experiences with smartphone technology, highlighting the importance of aligning the app design with the prevailing interface standards (Spann and Stewart, 2018).
4.3.4. M1. User involvement
The activeness level of users significantly influences strategies related to user participation, such as user support and controllability. In many human-computer interaction models, users are typically categorised as ”active process operators” or ”passive process operators” (Persson et al., 2001). This distinction is often observed in shared technology scenarios, such as health technologies utilised by physicians, nurses, and patients (Xu et al., 2014). Here, users may act as active users, with direct control over the technology, or passive users, who interact with the technology but lack direct control (Montague and Xu, 2012; Inbar and Tractinsky, 2009). In our research, we categorise active and passive users according to their willingness to interact with the system. This variation in involvement is not determined by shared technology but refers to the extent to which users actively or passively engage and comprehend the system (Eiband et al., 2021; Gómez-Carmona et al., 2024). Baroni et al. (2022) highlight the significant role of user involvement in the acceptance of intelligent systems. Recognising the level of user involvement is essential for customising system behaviour, as indicated by Oertel et al. (2020). User involvement levels vary depending on the context, task, and individual characteristics (Ramos et al., 2019). In our user study, we identified two types of user involvement. Active user involvement is exemplified by participants who proactively experiment with various data sources to understand how they impact the system’s output (FP 4,6,7, IP 3), desire for active participation and a sense of control over the adaptation process (FP 8,11). They are active in the adaptation process, willingly exploring and approving various adaptation suggestions (Gajos and Chauncey, 2017) (
). On the contrary, passive user involvement characterises individuals who are more inclined to seek information about how the adaptation works but do not actively provide feedback or corrections to the system (IP 2,4,6, FP 1,2,3,5,9) (
). Passive participants may stop exploring once they believe the current UI meets their minimum requirements (Eiband et al., 2021; Parra and Brusilovsky, 2015). Participants with a passive involvement with the app, with some admitting they ”have not fully explored what the app can do” (FP 9), others expressing ”a lack of concern about the system’s adaptation process” (IP 2), some indicating tolerance for most generated adaptations (FP 5), and some preferring minimal interaction with the app (FP 2). As Figure 5 denotes, participants can actively experiment with the way different settings for adaptations influence system output if the software provides switches and configurable options within the UI (
). For example, they can access a dashboard for the adaptation types and configure the settings for different adaptations. On the contrary, users with a negative perception of the software show less enthusiasm for this level of participation. They tend to prefer less or even no interaction when it comes to experimenting with adaptation settings (
Passive user involvement can also be influenced by other roles within the system, such as caregivers and family members (FP 8,9,10, IP 1,3,4,5,7). Participants highlighted the importance of involving family members or caregivers in the use of the system, as caregivers often play an active role in medical care decisions (Xie, 2009; Cassileth et al., 1980). Allowing others to handle the adaptation process may alleviate cognitive effort for older users (Schiavo et al., 2022), but could potentially reduce their independent use of the technology (Munteanu and Salah, 2017). This multifaceted spectrum of user involvement highlights the diverse ways in which users engage in the adaptation process, necessitating a tailored approach to accommodate these varying levels of involvement.
4.3.5. M2. User experience with mHealth applications.
Previous experience with technology plays a crucial role in shaping users’ perceptions and interactions with such systems (Parameswaran et al., 2015). This previous experience influences users’ perceptions of effort expectancy, which in turn affects their intention to use certain technologies (Venkatesh et al., 2012). In the realm of health-related applications, research has shown that users are more inclined to adopt mHealth applications when they perceive them to be easy to use, resulting in lower effort expectancy (Sun et al., 2013). In our user study, the experience with mHealth was measured by how frequently they use health apps since the standard to enter our user study is having the experience of using the mHealth application. Our user study highlights the impact of past experience with mHealth applications on users’ propensity to control adaptations within the system. Experienced users are more inclined to explore control options for adaptations, possibly because their prior experience with the system reduces their cognitive load, encouraging them to explore new actions and functionalities (Albers, 1997). Based on the insights in Figure 5, users who have experience with mHealth apps tend to advocate for greater control over adaptations, often suggesting detailed designs for these adaptations. Their familiarity with mHealth applications influences their preferences about how adaptations should function, as they seek personalised ways to integrate these adaptations into the system (
). Participants with less experience with mHealth technologies often struggle to conceptualise how adaptations should occur. Research indicates that experience can enhance the accessibility of knowledge in memory (Fazio and Zanna, 1978). Therefore, less experienced users typically expect the system to provide sufficient assistance with minimal cognitive burden when implementing adaptations (
). This is consistent with findings suggesting that individuals with less experience with a specific technology may value technical assistance more (Taylor and Todd, 1995). Limited familiarity and knowledge of the technology can lead to greater reliance on support. It should be noted that several participants expressed concerns during user sessions regarding the potential impact of adaptations on their technologically inexperienced parents (IP 1,5,7). This demographic, marked by limited technological know-how, faces challenges in navigating any form of technology and may feel hesitant to explore adaptations for fear of making mistakes while interacting with the applications.
4.3.6. M3. Health conditions
Chronic diseases exert varying physical, psychological, and mental impacts on participants (Harvey et al., 2012; di San Marco et al., 2019; Lorig and Holman, 2003; WHO, 2022). Our findings suggest that extensive control and support for adaptation may not be suitable for individuals lacking the capacity to make such decisions, especially if they are dealing with many lifestyle and treatment choices (Scherer et al., 2021). This observation is of particular relevance for individuals with severe health conditions or those who have recently been diagnosed with chronic diseases (Pizzoli et al., 2019). This aligns with the findings of Zhang et al. (2021), who noted that participants with severe mental health conditions faced challenges in customising certain messages within the app, primarily due to their emotional and cognitive states. Certain participants who consider themselves experts in managing their health conditions tend to become their own primary caregivers and demonstrate a greater willingness to explore various adaptation possibilities. As shown in Figure 5, individuals experiencing more severe symptoms and facing greater challenges in managing their health express a stronger preference for simplicity and greater system assistance. For example, users with more severe symptoms, advanced age, and multiple chronic diseases (FP 9, IP 4,5) tend to prefer adaptations occurring in specific parts of the interface that are easy to predict and understand (
). Additionally, they expect more guidance, such as instructions and on-boarding assistance, for adaptations to alleviate their initial login or troubleshooting burdens. It is crucial to recognise the impact of comorbidity and the prevalence of multimorbidity in chronic diseases on challenges and coping mechanisms (Islam et al., 2014). Users with healthier conditions are more inclined to thoroughly explore the entire system, seeking to engage with various adaptation options. Their proactive approach stems from their health status, enabling them to experiment with different features and functionalities within the application(
4.4. Perspectives on our AUI-based prototype
To gather diverse insights, we supplemented our qualitative analysis with a post-session survey administered to all participants after the interviews and focus group sessions. This survey covered various aspects, including perceptions of the three main types of adaptations and preferences for different adaptation aspects discussed during the focus group or interview session, and any additional open-ended suggestions. In total, 22 participants in our user study responded to our post-session survey.
4.4.1. Opinions towards different adaptations
The post-session survey began by asking participants about their preferences for the different adaptations they encountered during the user study. Figure 6 illustrates the views of the participants about their preference for the three main types of adaptations. Content adaptation emerged as the most popular adaptation among participants, with 50% expressing a strong preference for it. This inclination is supported by users’ expressed interest in content adaptation during the focus group and interview sessions. A considerable number of responses questioned the benefits of certain presentation adaptations, particularly by younger adults (e.g., FP 8 and IP 6), thereby reflecting participants’ preference for specific types of presentation adaptation. Only 18% and 23% of the participants demonstrated a strong preference for ”Color and theme” and ”Font size” adaptations, respectively. A smaller group of participants raised concerns about the utilisation of specific avatars and animations in adaptations during focus group and interview sessions. Nevertheless, 36% of the participants showed a strong preference for their incorporation in adaptations.
”Change text to image” and ”easy-to-understand languages” are strongly favoured by more than half of the participants. This aligns with the information collected during the previous sessions, where the participants expressed appreciation for adaptations that simplified the language and incorporated visual elements. Many participants shared their challenges in understanding certain medical content, despite their higher education levels, highlighting the importance of accessibility in health-related information (e.g., FP 4 and IP 5). ”Hide the elements” and ”reorder the elements” emerged as the less popular types of content adaptation among the participants. This sentiment was echoed during the prior sessions, where participants cited issues with hidden elements they needed or reordered the content in a way they found undesirable (4.2.1). The issue was exacerbated by examples in the AUI prototype, where hiding and reordering behaviours were primarily controlled by the system. Participants may feel more comfortable if they have full control over what elements are hidden and how content is reordered.
The preference for behaviour adaptation was notably lower among participants, and only 32% expressed a more than neutral opinion towards it. Participants sought help to understand the intricacies of behaviour adaptation in the prior session, which inherently involves multiple steps, making it challenging to grasp. Several participants (e.g., FP 4,6,9) even commented that specific types of behaviour adaptation were not helpful, and some adaptations, such as pervasive strategies, were deemed ineffective in motivating them to use the application, especially if the application itself lacked attractiveness. This deviates from their initial expectations of adapting content to make it easier to use, potentially affecting the actual benefits or utility of the functions and the overall use of the applications. However, multi-modal interaction and navigation adaptation are favoured by most participants. This demonstrates how users connect with functionalities that enhance their lives, emphasising the crucial role of technology in meeting user requirements (Trajkova and Martin-Hammond, 2020).
4.4.2. Opinions towards different types of data
A significant proportion of study participants emphasised the importance of motivation and feedback as primary data sources for adaptation. In particular, participants emphasised in the prior sessions that they wished the system to evolve based on their feedback, particularly concerning adaptations they deemed less useful or were hesitant to adopt. Participants expressed significant concern about their social activity and emotional data compared to other types of data, and approximately 40% of the participants expressed a slight or no preference for these types of data. This may be due to privacy concerns about these two types of data, which were emphasised by younger participants in interview sessions. Smartphones and wearable sensors emerged as the most popular and preferred methods for data collection, reflecting the challenges participants faced with manual input. Participants have adopted strategies to streamline data entry into the application, such as scheduling data input times at midnight to minimise interruptions or utilising calendars to track blood sugar levels (FP 7,9).
4.4.3. Opinions towards automation of adaptation
Users were surveyed about their preferences regarding control of the system within the application, with options ranging from fully manual to fully automatic. The majority (73%) expressed a strong preference for a semi-automatic system, where the burden of adaptation is shared between the user and the system. Concerns are raised regarding both fully automatic and manual systems. Manual systems face criticism due to concerns about inadvertently altering settings or data, prompting users to seek assistance (e.g., FP 4,6,9). An automated system can also cause anxiety due to unforeseen changes made without prior notice.
4.5. Mapping of challenges to strategies
Drawing from our data analysis and the memos recorded while following the STGT method, we have found some insights spanning various categories. Figure 7 illustrates the correlation between the challenges identified (Section 4.2) and the strategies outlined (Section 4.3). This mapping facilitates a comprehensive understanding of the relationship between specific challenges and the corresponding strategies.
4.5.1. Trade-off between user burden, user support and controllability
( , , ) Users have consistently highlighted the importance of having various options for controlling and supporting adaptations. However, effectively managing the complexity of providing users with the appropriate level of control and support remains a challenge that can affect the usability and effectiveness of AUIs (Paymans et al., 2004; Bunt et al., 2004; Eiband et al., 2021). User control over adaptation can yield significant benefits by empowering users to tailor their experiences to better suit their needs and preferences (Parra and Brusilovsky, 2015; Kay, 2001). Additionally, user-controlled adaptation fosters a sense of ownership and agency, leading to increased engagement and motivation to utilise the system effectively (Sundar and Marathe, 2010; Sundar, 2008). Our study participants expressed a preference for control mechanisms such as dedicated adaptation dashboards, selective system component adaptation, and configurable adaptation factors. However, our findings also highlight the importance of considering users’ competence levels, the potential cognitive load, and the risk of adaptation processes becoming intrusive (Kurtz, 2005). Allowing excessive control may lead to distraction and inefficiency, particularly among users who lack the necessary knowledge or interest to make informed decisions (Jameson and Schwarzkopf, 2002) ( ). Furthermore, as discussed by Kay (2001), user preferences for control can vary significantly. Some users may feel obligated to have control over the adaptation and other elements of the user interface, while others may be less inclined towards control (Bunt et al., 2004). Providing users with an excessive number of control options can result in distraction and fatigue when making decisions ( ). Additionally, differences in user perspectives on controllability may also be related to specific tasks user performs, as highlighted by Gajos and Chauncey (2017). This relationship, focusing on task-related characteristics, is further explored in the study by Lavie and Meyer (2010). They found that the proportion of routine and non-routine tasks, as well as the efforts required for certain tasks, can affect the actual performance of AUIs. This underscores the importance of considering these factors, as users often engage in a mix of tasks when using mHealth apps, with some tasks being more frequent or demanding than others (Bunt et al., 2004). In tasks that are essential and frequently performed, the importance of user controllability diminishes, and users may prefer the system to handle these tasks without requiring explicit confirmation or manual adjustments (Peissner and Edlin-White, 2013). Conversely, for tasks that require more effort, users might expect to have greater control over the adaptation. Therefore, delegating control to users requires considering various factors, including the nature of the task and user characteristics, rather than focusing solely on one specific aspect.
Apart from the controllability, research has attempted to support the adaptation process. For example, some studies have explored the use of animated transitions to demonstrate the adaptation process to users (Dessart et al., 2011), while others have investigated the effectiveness of providing detailed explanations (Kühme, 1993). However, it seems questionable whether these methods should be applied to adaptations that attempt to compensate for users’ perceptual, cognitive, and motor impairments. The provision of support materials may not necessarily reduce cognitive demands, as additional support itself can impose an additional burden on users (Kulesza et al., 2013). Meanwhile, users may not always find value in support materials, such as explanations, especially when they lack control over the adaptation process, perceive that the system is functioning effectively, or consider the time and effort required to understand the information outweighing the benefits (Bunt et al., 2012). Therefore, while enhancements in controllability and support are beneficial, it is crucial to carefully consider their implementation and ensure that they meet users’ actual needs ( ).
4.5.2. Trade-off for usability issues
( , , ) AUIs are increasingly seen as a solution to cope with the growing diversity of usage contexts, devices, and users (Browne et al., 1990; Höök, 1998; Norcio and Stanley, 1989). They offer solutions to various usability issues in mobile applications, including improving accuracy, efficiency, and user learning, as well as addressing information overload and helping in the use of complex systems (Browne et al., 1990; Höök, 1998). However, previous research on AUIs has also demonstrated a trade-off between adaptive mechanisms and usability (Paymans et al., 2004; Jameson, 2007; Mezhoudi et al., 2015; Deuschel, 2018; Höök, 2000; Glass et al., 2008; Bunt et al., 2004; Gajos et al., 2006, 2008; Turk et al., 2019). We found several usability challenges associated with AUIs, including privacy concerns, predictability issues, comprehensibility difficulties, and UI obtrusiveness in our user study. Interestingly, our study reveals that certain strategies proposed by participants to enhance the adaptation design can alleviate or exacerbate some existing usability challenges. For example, aligning visual elements and icons with user preferences can reduce UI obtrusiveness ( ). However, it is crucial to note that introducing icons like lifelike characters, such as a doctor icon or an ambulance icon, may create unrealistically high expectations regarding the system’s competence in understanding and adapting to the user (Jameson, 2007; Sproull et al., 1996). Creating the right expectations in users is crucial for effective design, ensuring neither overestimation nor underestimation of the system’s competence (Höök, 2000). Mismatches in expectations can directly reduce trust and subsequently lead to a decreased intention to use the interface again (Glass et al., 2008; Pu and Chen, 2006). Other strategies, such as attempts to enhance control or privacy over the adaptation, may necessitate additional user interactions for proposed adaptations, data inputs to extract feedback for the generated adaptation, or system notifications for changes made through adaptations. This could potentially increase the risk of disruption and divert user attention from their previous focus (Jameson, 2007; Deuschel and Scully, 2016) ( ).
We have observed that while certain strategies can mitigate specific usability challenges, they may inadvertently exacerbate others. Meanwhile, the effectiveness of a strategy depends on its appropriate application, as it can have varying impacts on usability. Despite the initial advantages of a strategy aimed at addressing a specific usability challenge, its excessive application can lead to the negation of potential benefits and the emergence of new usability issues (Jameson, 2007). For example, aligning with chronic disease guidelines and providing explanatory materials can improve adaptation comprehensibility ( , ). However, an excessive provision of support content can become obtrusive, difficult to grasp, and impair user learnability (Eiband et al., 2021; Paymans et al., 2004). Similarly, step-by-step adaptation at regular intervals could improve UI predictability and could also increase the intrusiveness of the adaptation process ( ). However, constantly adapting the UI can be obtrusive and disrupt the workflow of users (Wesson et al., 2010). Future research efforts should aim to improve the benefits offered by AUIs while mitigating potential challenges associated with them.
Mitigating specific usability challenges in AUI design often entails trade-offs with other usability goals, emphasising the intricate nature of achieving a balance between adaptation and usability, and necessitating efforts to minimise conflicting demands. The key lies in striking a balance between the various demands of an interface. Our findings highlight two key considerations to deal with conflicting demands.
1) User priorities. Users often have varying priorities regarding conflicting usability goals (Rich and McGee, 2004). For example, in our user study, we observed that some participants preferred interfaces with consistent layouts and design elements across various screens, as it fosters familiarity and ease of use. In contrast, others prioritise adaptability and enjoy exploring new interface designs and features, even if it means sacrificing some consistency. Similarly, users may have different privacy priorities. While some individuals prioritise privacy and are cautious about sharing personal information, others are more inclined to offer data in return for customised user interfaces. These examples highlight the critical importance of offering alternative solutions that accommodate the diverse usability priorities of users (Bunt et al., 2004; Gajos et al., 2008). This approach assumes that different users may be open to trade-offs among different usability aspects. Harper et al. (2015) found that users reported usability issues with a system but remained overall satisfied because the benefits of having control options outweighed the encountered inconveniences.
2) Granularity. Usability objectives such as predictability, comprehensibility and controllability can be achieved at different levels of granularity (Bunt et al., 2004; Jameson, 2007). Our observations revealed that making minor adjustments during the adaptation process could not significantly impact the overall user experience. Instead, small incremental changes could alter the user’s familiarity with the UI, leading to a less consistent and less predictable user experience (Gajos et al., 2006; Bunt et al., 2004). Additionally, these low-level granularity adaptations may have a limited impact on the comprehensibility of the adaptations, as they typically affect only small and limited parts of the UI (Abrahão et al., 2021). High-level granularity adaptation involves modifying multiple aspects of the UI simultaneously (Cristea and Calvi, 2003; Zeidler et al., 2013), such as altering the overall colour scheme, font styles, navigation menus, and placement of key features at the same time. It may introduce more usability challenges, as it could be perceived as a completely new app from the user’s perspective. Typically, high-level granularity adaptations occur less frequently and may occur only once to tailor the app to the user’s specific needs, which could compensate for some usability issues. This approach is particularly advantageous when users first engage with the system, as they do not have preconceived expectations of its appearance or functionality (Bunt et al., 2004). In such scenarios, users approach the system without investing time or effort into learning its interface or forming expectations about its design and functionality. Therefore, any adaptation that occurs may not significantly affect the predictability or comprehensibility of the UI, as users lack a pre-established reference point for comparison. However, it may pose challenges for users who are less familiar with such applications and lack insights into which design and functionality would best suit their needs (Gajos et al., 2006).
Finding the right balance between these factors is essential for designing AUIs that effectively manage the trade-off between adaptation and usability. Some usability issues introduced by AUIs can be mitigated without necessarily improving the user’s mental model of the adaptive system (Paymans et al., 2004; Rosman et al., 2014). However, systems that actively help users overcome usability issues tend to foster a deeper understanding and trust in the system’s ability to adapt effectively to their needs, encouraging sustained engagement and usage of mobile health apps (Martin-Hammond et al., 2019). As users become more familiar with the adaptive capabilities and benefits of the system, they are likely to develop a deeper appreciation of its role in enhancing their overall experience and well-being.
4.5.3. Trade-off between independence and assistance
( , ) User support typically focuses on facilitating the execution of app functions, but this type of support may hinder user learning process and diminish their overall experience with the system (Lanier, 1995). Learning involves acquiring skills, knowledge, and competencies within a specific domain, which allows individuals to evolve and become more independent. However, this learning process often entails significant costs, including the time and effort spent acquiring and executing activities, which can be error-prone and time-consuming ( ). It can pose a significant challenge for users when they are required to learn to use the system each time it undergoes adaptation. In addition to providing user support, mitigating the cognitive effort required to adapt the system can be achieved by allowing other users, such as caregivers or health professionals, to manage the adaptation process. For example, a chronic disease management app could be tailored by caregivers or health professionals, albeit at the expense of the user’s independent utilisation of the technology ( ). The decision to invest time in domain-specific learning or delegate tasks to the system should ultimately be made by the user. Offering users a spectrum of options between independent and assisted app usage may still be beneficial. The design of support features should align with users’ objectives and contextual factors (Fischer, 2006), accommodating users’ preferences and comfort levels in embracing challenges across various situations.
The inevitability of trade-offs complicates decision-making in UI adaptation, as users frequently differ in their prioritisation of conflicting goals, level of competence, objectives for using the app, and the specific usage scenario. It is imperative to provide alternative solutions that cater to users with varying priorities (Billsus and Pazzani, 2007). Furthermore, exploring innovative UI design solutions can help mitigate the need for users to choose suboptimal options. For example, Cheverst et al. (2005) developed an office control system prototype that initially offered users two task delegation options. Despite these choices, users were dissatisfied with the usability of the prototype. However, acceptance markedly improved when designers expanded the design space by introducing a separate small screen for the office control system.
5. Findings from survey
5.1. Survey participants
Concurrently with our focus groups and interviews, we surveyed 90 participants with chronic diseases, all of whom have experience using mHealth applications. Their detailed demographic information is summarised in Table 4. Most of the participants identified as men (56%), between the ages of 18 and 74. In particular, the age group of 25-34 constituted the largest group, comprising about 33% of the total participants. There was representation from older age groups, with 16% of participants over 55 years of age and 21% between 35 and 44 years of age. Geographically, our survey captured responses from a diverse set of countries, with approximately half of the responses originating from Australia. China and the USA are also significant contributors, accounting for 23% and 13% of the responses, respectively. In terms of educational background, a bachelor’s degree is the most common attainment (44%), followed by 34% with education levels below a bachelor’s degree and 21% with higher degrees, such as master’s or doctorate degrees.
We grouped participant chronic health conditions into four broader groups: cardiometabolic (e.g., diabetes, high blood pressure, obesity, and heart disease), respiratory (e.g. allergies, asthma, and chronic lung disease), immune-related (e.g., cancer, Parkinson’s disease, and compromised immune system), and mental health conditions (Camacho-Rivera et al., 2020). Cardiometabolic diseases were the most prevalent, reported by 52% of the participants. It is worth noting that some participants reported multiple chronic diseases, such as having both diabetes and high blood pressure, or a combination of cardiometabolic disease and another type of disease, such as asthma (a respiratory disease). Indicates a phenomenon known as multimorbidity (Islam et al., 2014), where individuals may experience two or more chronic conditions simultaneously. This is also consistent with other studies (Anderson et al., 2016). Participants were asked to indicate their familiarity with mHealth apps. As shown in Table 4, respondents reported using various types of mHealth apps, with health-promoting and self-monitoring apps being the most prevalent, utilised by 84% of participants. These applications typically cater to functions related to fitness, medication, and diet, which is consistent with findings from previous research (Kayyali et al., 2017). This prevalence can be attributed to the ongoing need for individuals with chronic diseases to monitor their health status and adhere to prescribed treatment regimens (Hamine et al., 2015). The primary motivations cited for using mHealth apps include symptom monitoring (58%), promoting physical activity (54%), and managing dietary intake (47%). Participants were also asked about the frequency of using the mHealth app (see Figure 8). In our survey study, 54% of participants reported using health apps daily, while 33% reported weekly usage. Most users spend between 1 to 20 minutes per session, aligning with findings from prior research (Robbins et al., 2017; Kc et al., 2021). Users who access the app two or more times per day typically allocate a shorter duration per session. On the contrary, people who use the app less frequently (less than once a month or less than once a week) tend to engage in longer sessions. The former group may use the app for quick, frequent interactions or brief check-ins throughout the day, while the latter may delve into more comprehensive interactions during less frequent usage, possibly for specific tasks or content consumption.
| Demographics | # | % of Participants |
| Age of the Participants1 | ||
| 18-24, 17/19% — 25-34, 30/33% — 35-44, 19/21% — 45-54, 10/11% — 55-64, 11/12% — 65-74, 3/3% | ||
| Gender of the Participants2 | ||
| Female, 38/42%, — Male, 50/56%, — Prefer not to say, 2/3% | ||
| Countries of the Participants | ||
| Australia | 44 | 49% |
| China | 21 | 23% |
| USA | 12 | 13% |
| UK | 6 | 7% |
| Other (Nigeria, Canada, Korea, Spain and Sri Lanka) | 7 | 8% |
| Education information of the Participants3 | ||
| Less than Bachelor’s degree | 31 | 34% |
| Bachelor’s degree | 40 | 44% |
| Postgraduate (Master’s and Doctoral degree) | 19 | 21% |
| Categories of chronic disease4 | ||
| Cardiometabolic | 47 | 52% |
| Immune-related | 31 | 34% |
| Mental health conditions | 7 | 8% |
| Respiratory | 11 | 12% |
| Type of mHealth application used 5 | ||
| Health promoting and self-monitoring application | 76 | 84% |
| Informative application | 41 | 46% |
| Assistive application | 29 | 32% |
| Communication application | 24 | 27% |
| Health game application | 16 | 18% |
| Rehabilitation application | 16 | 18% |
| Exercise application | 7 | 8% |
| Main reason to use the app | ||
| To monitor my chronic disease symptoms | 52 | 58% |
| Increase my physicial activity levels | 49 | 54% |
| To track what I eat | 42 | 47% |
| To get education about my chronic disease | 34 | 38% |
| Help with weight loss | 28 | 32% |
| Manage my medications | 26 | 29% |
| 1,2,3: The percentage does not strictly add up to 100% due to rounding. 4: The percentage does not strictly add up to 100% due to multimorbidity. 5: The percentage does not strictly add up to 100% due to the use of multiple applications. | ||
| Aspects of adaptations | # | % of Participants | Aspects of adaptations | # | % of Participants |
| Different types of adaptation | Data source of adaptation | ||||
| P: Graphic design | 52 | 58% | UC: Physiological characteristics | 53 | 59% |
| P: Information architecture | 30 | 33% | UC: Physical characteristics | 52 | 58% |
| P: Sound effect | 25 | 28% | UC: Preference | 48 | 53% |
| C: Content complexity | 53 | 59% | UC: Psychological characteristics | 47 | 52% |
| C: Interface elements rearrangement | 42 | 47% | UC: Demographics | 37 | 41% |
| B: Multimodal interaction | 41 | 46% | UC: Social activity | 29 | 32% |
| B: Difficulty level | 39 | 43% | IR: Feedback | 41 | 46% |
| B: Add on functions | 36 | 40% | IR: Interaction with the interface | 34 | 38% |
| B: Navigation adaptation | 31 | 34% | IR: Emotions | 34 | 38% |
| B: Different persuasive strategy | 31 | 34% | IR: Performance in game | 29 | 32% |
| TS: Goals | 40 | 44% | |||
| P: Presentation adaptation, C: Content presentation, B: Behaviour adaptation | TS: Motivation | 35 | 39% | ||
| UC: User characteristics, IR: Interaction related, TS: Task specific | TS: Role | 24 | 27% | ||
| Different data collection method | Preferred level of involvement in the adaptation | ||||
| Smartphone sensor | 61 | 68% | Semi-automatic | 49 | 54% |
| Wearable sensor | 58 | 64% | Automatic | 34 | 38% |
| User input through the application | 49 | 54% | Manual | 7 | 8% |
| Analysis of user behaviour through the application | 41 | 46% | |||
| Analysis of activities with keyboard | 27 | 30% | |||
| Variables | Categories | CC* | AD* | DP* | MI* |
| Age group | 45-74 | 3.212/(0.065) | 3.764/(0.02) | 0.637/(0.43) | 5.824/(0.002) |
| Nationality | UK | 0.066/(0.16) | 0.825/(0.843) | 0.889/(0.908) | 0.133/(0.095) |
| USA | 0.276/(0.107) | 0.299/(0.178) | 1.504/(0.567) | 0.346/(0.193) | |
| China | 0.125/(0.003) | 0.215/(0.025) | 1.179/(0.796) | 1.081/(0.896) | |
| Other | 0.783/(0.803) | 0.341/(0.264) | 1.2/(0.845) | 0.533/(0.492) | |
| Chronic diseases | Mental health | 4.173/(0.033) | 0.358/(0.093) | 2.126/(0.181) | 0.661/(0.475) |
| Cardiometabolic | 12.215/(0.013) | 8.093/(0.029) | 11.264/(0.006) | 1.782/(0.471) | |
| Immune-related | 0.536/(0.554) | 3.823/(0.168) | 2.293/(0.339) | 1.463/(0.664) | |
| * Odds Ratios/(P-value),(CC=Content complexity, AD=Add on functions, DP=Different persuasive strategy, MI=Multimodal interaction) | |||||
5.2. Different types of adaptations
In our earlier SLR on AUIs for chronic disease support (Wang et al., 2023b), we compiled a comprehensive list of adaptations implemented in chronic disease-related applications by researchers. Our aim with this survey is to discern the users’ perspectives regarding the significance and value attributed to these adaptations. Our survey revealed a diversity of preferences among the respondents, with no singular adaptation type dominating over others. Content complexity (59%) and graphic design (58%) emerged as the most prevalent, closely followed by interface elements rearrangement (47%) and multimodal interaction (46%). In contrast, the sound effects exhibited the lowest frequency among the listed adaptations (28%) (see Table 5). This finding is not consistent with prevailing trends in existing AUI studies within the domain of chronic diseases (Wang et al., 2023b). Our SLR underscored graphic design (presentation adaptation) as the predominant adaptation type, which encompassed layout, font size, colour, and theme adjustments, while other adaptation types received comparatively less attention. However, our survey findings highlighted users’ significant emphasis on content adaptation and some types of behaviour adaptation. The difference in research trends and user preferences offers valuable information on the types of adaptation preferred by users. It underscores the importance of aligning researchers’ efforts with user preferences and priorities to ensure the effective design and implementation of adaptive user interfaces in chronic disease management applications.
5.2.1. Relationship between users’ preferences for adaptations and their demographic characteristics
We conducted a Chi-square independence test to investigate the influence of demographic variables on user preferences for adaptation types. The results indicate significant associations between age, nationality, chronic diseases, and preferences for adaptations such as content complexity, add-on functions, persuasive strategy, and multimodal interaction. Subsequently, a binary logistic regression analysis was employed to examine the relationships between various adaptations and significant demographic factors identified in the Chi-square test (see Table 6). It is important to note that statistical models such as binary logistic regression are built on certain assumptions about the data, including independent observations and no perfect multicollinearity (Harris, 2021), which are met in our dataset. Compared to participants from Australia, individuals from China are less likely to prefer adaptations related to content complexity (OR = 0.125) and add-on functions (OR = 0.215). Individuals from China may have different expectations or preferences regarding the complexity of content or additional functions in mHealth applications, leading to a lower likelihood of preferring content complexity adaptations compared to participants from Australia. Participants with cardiometabolic diseases showed a significantly higher inclination towards adaptations such as different persuasive strategies, additional functions, and content complexity compared to those with respiratory diseases. Similarly, participants with mental health conditions also demonstrated a notable preference for content complexity. These preferences may arise from the self-management nature of cardiometabolic and mental health conditions, which often necessitate individuals to actively monitor their health and follow treatment plans (Cruz-Ramos et al., 2022; Chow et al., 2016; Torous et al., 2015). Furthermore, participants over 45 years of age exhibited a higher tendency to seek multimodal interaction (OR = 5.824) and additional function (OR = 3.764) adaptations compared to their younger counterparts. This tendency among older users may stem from a preference to minimise interaction efforts and an increased need for additional assistance with app navigation (Trajkova and Martin-Hammond, 2020; Betts et al., 2019; Martin-Hammond et al., 2019). As individuals age, they may encounter challenges related to vision, dexterity or cognitive abilities, making features such as multimodal interaction and additional functions particularly appealing to facilitate their interaction with mHealth applications.
5.3. Data sources for adaptation
Our SLR identified various sources of data used as a basis for adaptation (Wang et al., 2023b). In our survey, we aimed to further explore users’ perspectives on the preferred types of data for adaptation. We found that participants exhibited various preferences about the source of adaptation data and almost half expressed the desire for the app to adapt based on physiological, physical, and psychological characteristics, user preferences, and user goals (see Table 5). The user characteristics data was the most popular compared to the interaction-related data and the task-specific data. This finding is consistent with existing research trends (Wang et al., 2023b). Interestingly, there was a notable preference for interaction-related and task-specific data in the prior research. This underscores the complexity of user preferences and the need for nuanced approaches to adaptation in user interfaces.
5.3.1. Relationship between users’ preferences for data sources for adaptations and their demographic characteristics
We used a Chi-square independence test to explore how demographic variables influence users’ preferences for data sources for adaptation. The findings revealed significant associations between nationality, chronic diseases, education level, and various types of data (see Table 7). Subsequently, we performed a binary logistic regression analysis to examine the relationships between different data sources and the significant results obtained from the Chi-square test. Participants from different countries exhibit varying preferences for data adaptation based on physiological, physical, preference, feedback, goals, and motivation factors. Given the relatively small number of participants from the UK and USA in our survey, it is essential to acknowledge the potential impact on the statistical power of our analysis. With a reduced sample size, there is a risk of diminishing the ability to detect genuine effects or associations accurately. Consequently, the observed relationships between nationality and data adaptation preferences may be less reliable, introducing uncertainty into our findings (Nemes et al., 2009). Individuals with higher levels of education exhibit a stronger inclination toward the adaptation of their motivation to use the application (Bachelor: OR= 7.835, Postgraduate: OR= 6.798). Individuals with higher levels of education often have a greater awareness of the benefits of using technology for health management and may also have a greater understanding of the importance of motivation in achieving health-related goals. Therefore, people with higher educational levels can exhibit a greater tendency to adapt their motivation to use the application due to their improved understanding of the role of motivation in achieving health outcomes (Woldaregay et al., 2018). Individuals with cardiometabolic diseases are more inclined to desire app adaptations that align with their goals (OR=13.948). This preference could be attributed to the management of cardiometabolic diseases, which often involves monitoring the levels of diet and physical activity and striving to achieve specific goals (Cruz-Ramos et al., 2022; Chow et al., 2016; Torous et al., 2015).
Our survey reveals that the predominant methods of data collection include smartphone sensors (68%) and wearable sensors (64%) (see Table 5). The existing literature shows a pronounced focus on wearable sensors over smartphone sensors in mHealth applications targeting chronic diseases (Wang et al., 2023b). With the widespread adoption of smartphones, an increasing number of users prefer to use their smartphones for data collection rather than rely on other devices. In particular, no significant correlations were found between the data collection method and demographic variables.
| Variable | Category | Physiological characteristics | Physical characteristics | Preference | Feedback | Goals | Motivation |
| Nationality | UK | 0.093/(0.053) | 0.042/(0.019) | 0.047/(0.018) | 1.618/(0.633) | 0.222/(0.899) | 0.727/(0.767) |
| USA | 0.121/(0.009) | 0.032/(0.002) | 0.138/(0.018) | 0.205/(0.07) | 0.11/(0.015) | 0.168/(0.052) | |
| China | 0.303/(0.066) | 0.485/(0.283) | 0.127/(0.003) | 0.216/(0.016) | 0.129/(0.005) | 0.048/(¡.001) | |
| other | 0.224/(0.092) | 0.736/(0.747) | 0.624/(0.647) | 0.565/(0.512) | 0.238/(0.122) | 0.101/(0.057) | |
| Chronic diseases | Mental health | 1.055/(0.929) | 0.833/(0.783) | 0.202/(0.013) | 0.471/(0.193) | 1.682/(0.412) | 0.844/(0.801) |
| Cardiometabolic | 136.518/(0.999) | 6.172/(0.179) | 0.885/(0.895) | 4.012/(0.138) | 13.948/(0.034) | 3.352/(0.208) | |
| Immune-related | 0.354/(0.269) | 0.057/(0.017) | 1.244/(0.835) | 0.477/(0.456) | 2.954/(0.287) | 0.167/(0.17) | |
| Education level | Bachelor’s degree | 1.023/(0.97) | 1.09/(0.896) | 0.952/(0.937) | 1.428/(0.544) | 1.1/(0.881) | 7.835/(0.006) |
| Postgraduate | 2.116/(0.355) | 1.568/(0.602) | 1.71/(0.513) | 1.511/(0.569) | 1.634/(0.52) | 6.798/(0.035) | |
| * Odds Ratios/(P-value) | |||||||
5.4. Preferred level of involvement in the adaptation
The SLR found varied ways of the involvement of the user in the adaptation process. In the survey, we discovered that participants generally favoured a semi-automatic adaptation approach (54%), which involves collaboration between the system and the end-users to achieve adaptation (Mukhiya et al., 2020; Abrahão et al., 2021). On the contrary, only a small minority of participants (8%) expressed a preference for a fully manual system, where users have complete control over the modification of specific UI elements to suit their needs (Akiki et al., 2014). The limited preference for fully manual systems suggests that users may find manual adaptation processes cumbersome or time-consuming, a sentiment that participants in the user study articulated and which was further corroborated by their feedback in the post-session survey. Fully manual systems may also require users to have a higher level of technical proficiency, which could be a barrier for some individuals, particularly those with limited technological literacy or cognitive abilities (Akiki et al., 2014). The SLR also indicated a lower preference for manual systems, which aligns with our survey findings. While significant research efforts have been dedicated to automatic systems, there has been a recent surge in interest in semi-automatic approaches (Wang et al., 2023b). No significant correlations are found between the level of involvement and demographic variables.
5.5. Key findings from survey
Our survey offers valuable insights into users’ preferences regarding various aspects of design adaptations, which is partially consistent with existing research efforts. From our analysis, we have identified four key findings that help unravel the intricacies of user preferences. The survey results complement and provide context for the qualitative findings obtained from our interview and focus group studies.
Key findings 1: Multimorbidity nature of chronic diseases. The prevalence of multimorbidity among participants underlines the common experience of individuals managing multiple chronic diseases simultaneously. This phenomenon is not unique to our study and has been consistently reported and emphasised in other research (Islam et al., 2014; Anderson et al., 2016). However, despite the widespread occurrence of multimorbidity, much of the existing exploration in the realm of mHealth applications for chronic disease management has predominantly focused on addressing a single chronic disease (e.g., hypertension (Hallberg et al., 2016) and asthma (Tinschert et al., 2017)). Similarly, in the context of AUIs, research has primarily concentrated on investigating their efficacy in managing single chronic diseases such as stroke (Burke et al., 2009) and diabetes (Pagiatakis et al., 2020). Consequently, there is a lack of approaches that address the diverse and complex requirements of individuals facing the challenges of managing multiple chronic conditions simultaneously. Closing this gap is essential to ensure that mHealth interventions and AUIs effectively support individuals in navigating the complexities associated with multimorbidity. This phenomenon links to what could be adapted in the UI and users’ preference for controllability. Users may feel overwhelmed by managing multiple diseases, leading to a lack of motivation and interest in actively intervening or taking responsibility themselves (see Figure 9).
Key findings 2: Variations in purposes and usage pattern. The extensive use of mHealth applications for various purposes, ranging from symptom monitoring to dietary management, highlights their vital role in supporting the self-management of individuals with chronic diseases. Each app function serves a specific need, leading to different usage patterns among users. However, even within the same functionality, users may exhibit fluctuating usage patterns, with no fixed frequency or duration. Research indicates that individuals engage with physical activity tracking apps in intermittent intervals, characterised by periods of consistent use followed by breaks and subsequent reengagement (Lin et al., 2018; Meyer et al., 2016). These task-related characteristics have implications for the design of AUIs and their performance, especially concerning users’ preferences for controllability over adaptations and the timing and frequency of adaptation (Lavie and Meyer, 2010; Peissner and Edlin-White, 2013) (see Figure 9). For example, tasks that require minimal effort and are performed daily, such as recording mood, stress levels, or blood pressure, may not require manual control by the users. In such cases, users might prefer the system to handle these tasks automatically without requiring explicit confirmation. Additionally, since these tasks are performed regularly compared to other functionalities, users may not desire frequent changes to occur each time they log in. Detailed discussions on task characteristics and adaptation strategies can be found in Section 4.5.1.
Key findings 3: Cultural and contextual differences. The analysis identified significant correlations between demographic variables such as age, nationality, and chronic diseases, and preferences for specific types of adaptations. For example, Chinese individuals exhibited a lower inclination toward content complexity adaptations compared to Australian participants, suggesting cultural and contextual differences in adaptation preferences. Moreover, individuals with higher levels of education demonstrated a stronger tendency to adapt their motivation to use the application, likely due to their deeper understanding of motivation’s role in health outcomes. Previous research highlights the importance of culturally specific design preferences in influencing system usefulness across different cultural contexts (Evers and Day, 1997). Various cultural factors significantly influence individuals’ preferences regarding UI perceptions and design. For example, religion plays a role in shaping users’ preferences for specific symbols and colours, influencing their implicit trust in the interface (Siala et al., 2004). Political orientation and social structure may affect users’ perception of the hierarchy and complexity of information presentation (Schmid-Isler, 2000). Individuals with different levels of education may require varying levels of support and cues from systems (Reinecke and Bernstein, 2013). When considering the actual adaptation process, cultural and contextual factors can significantly influence whether the adaptation meets the user’s expectations or not, potentially resulting in a mismatch with their cultural background. Additionally, user education levels can play a role in determining the ideal timing to initiate adaptation and the extent of support needed (see Figure 9).
Key findings 4: Preference aligns with goal. Through our survey analysis, we discovered that different types of chronic disease can significantly influence the adaptations and data preferences of users. Specifically, participants with cardiometabolic diseases showed a greater inclination towards app adaptations that align with their usage goals. This preference likely arises from the pressing need for self-management inherent in cardiometabolic diseases. The availability of various applications for the management of chronic diseases or multimorbidities in app stores underscores the importance of adaptation aligning with users’ primary app usage goals and adding value to their overall app experience (Bricca et al., 2022). This resonates with the overarching categoryobserved in the user study section (see Figure 9), emphasising the essential value brought by adaptations and their alignment with users’ primary app usage goals (Trajkova and Martin-Hammond, 2020). Understanding these dynamics is crucial for developers and designers aiming to create adaptations for mHealth applications that not only support individuals with chronic diseases but are also valued by them. There are numerous options and elements available for adaptation on the UI, but it is essential to prioritise those that align with the user’s goals. However, selecting the perfect solution from such diverse options is challenging, as it depends on individual user experiences and the type of application and its adaptations, as suggested by the study conducted by Jameson and Schwarzkopf (2002). However, ongoing research continues to try to offer recommendations and guidance to users, helping them select the most appropriate options based on the data available within the system (Reinecke and Bernstein, 2013).
6. Guidelines for AUI Design for mHealth applications targeting chronic disease
Our approach to developing our guidelines follows and adapts the framework proposed by Hermawati and Lawson (2015) and Quiñones et al. (2018). The methodology proposed by Quiñones et al. (2018) outlines various stages that facilitate the development process and allow the formal specification of guidelines through appropriate validations and refinements. However, some of these stages lack detailed explanations of their implementation. The involvement of users in the process and the expansion of guidelines within a particular domain are not sufficiently emphasised. On the other hand, Hermawati and Lawson (2015) have modified a user-centric approach to create usability heuristics tailored to specific domains. Consequently, we integrate elements from both methodologies into our guideline development process. The guideline developmrnt process is as follows (see Figure 10):
Exploratory stage: During our review of the literature, we gathered essential information to develop a new set of guidelines for designing AUIs in mHealth applications. This involved examining existing guidelines, design considerations, and recommendations pertinent to our research focus. An analysis of our literature review findings is presented in Section 6.1.
Experimental stage: We have conducted a user study comprising interviews, focus groups, and a survey to delve into users’ experiences, preferences, and challenges related to adaptive user interfaces in mHealth applications. Through the analysis and interpretation of these qualitative and quantitative data, we obtained valuable information to inform the creation of guidelines.
Descriptive stage: In the formalisation process of the main guidelines, we began by selecting and highlighting the most significant categories and concepts derived from all qualitative and quantitative data collected during the experimental stage. Initially, we grouped the needs identified from users and then aligned this information with the guidelines and design recommendations uncovered during the exploratory stage (Hermawati and Lawson, 2015). For any remaining information that cannot be directly mapped, we abstract new guidelines. Furthermore, we provide detailed information pertaining to the application domain, including definitions and illustrative examples to enhance clarity and applicability. The refined and finalised guidelines are available in Table 10.
Validation stage: Given the critical nature of this stage, which directly pertains to the specific domain, we prioritise the solicitation of feedback from the users we have in the interview and the focus group study, as recommended in the literature (Hermawati and Lawson, 2015). Further details regarding the validation process can be found in Section 6.2.
6.1. Literature review over existing guidelines or design recommendations
[…]
| Study | Target audience | Target domain | Design considerations/recommendations/criteria |
[…]
| Guideline | Definition | Purpose | Example |
[…]
| (EASY) G1: Allowing Controlling and Changing Things | How much users can control and change things during adaptation. | A good mix between letting users control things and having the system do some stuff automatically. | |
[…]
| (EASY) G2: Helping Users Easily Use Adaptive Interfaces | Help users to use adaptive interfaces more easily. | Make sure users know how adaptive features work, what they can do, and how to use them properly. | |
| G3: Aligning Adaptations with User’s Chronic Disease | Consider multimorbidity, the understanding, and the progression of their chronic diseases. | Ensure that any adaptations made do not hinder but rather improve users’ ability to effectively manage their chronic condition. | ”Including relevant icons and visuals related to chronic diseases.” (IP 1) ”The process of adaptation needs to consider various chronic diseases, since individuals with certain types of disease may have varied preferences for adaptation.” (From the survey data analysis) |
| (EASY) G3: Considering Users’ Chronic disease | Look at if users have multiple health issues, how well they know their conditions, and how they handle their health problems. | Make sure that any changes do not make it harder but instead help users better manage their health condition. | |
| G4: Aligning Adaptations with App Usage Patterns | Ensure that adaptations align with the frequency, duration, and effort of usage for specific functions. | Improve the effectiveness of implemented adaptations, enhance user experience, and engage users without disrupting their typical app usage. | ”Tasks that require minimal effort and are performed daily may not require manual control from users.” (From memos) ”Users may not desire frequent adaptation to occur if it is a function they will use everyday” (From memos) |
| (EASY) G4: Following the App Use behaviour | Making sure that adaptations match how often, how long, and how much effort users put into using specific functions. | Create a smooth and engaged user experience by having the adaptations. | |
| G5: Aligning Adaptations with User’s Coping Style | Ensure that adaptations are in line with the coping strategies users use to deal with stressful situations. | Adapt the adaptations based on varying health statuses can enhance user engagement with the mHealth application and improve self-management of chronic disease. | ”I prefer frequent and clear reminders for any negative measurement. I do not see the point in avoiding the bad measurement.” (IP 5) ”I haven’t been using the app recently because I have not experienced any negative effects. I prefer not to see my high blood pressure measurement daily because I am afraid of seeing the bad measurement.” (IP 4) |
| (EASY) G5: Matching how users handle stress | Make sure that the changes made to the app match how users handle stress that casued by different health situations. | Tailor the features according to different health conditions can make the mHealth app more engaging and help users better manage their chronic diseases. | |
| G6: Collaborative Adaptation between Users and Caregivers | A cooperative approach in which both primary users and their caregivers work together to customise and fine-tune various aspects of the application, or even jointly utilise the application. | Emphasise the significance of the caregiver’s role throughout the adaptation process and explore strategies for integrating the caregiver’s contributions into the adaptation procedures. | ” I have to look over the health data, understand the app features, and set up some reminders in the app for my family members.” (IP 1,3,4, FP 9) ”I have to set up the app for my parents if the app has provided some flexibility in the functionality or other aspects.” (IP 4) |
| (EASY) G6: Collaboration Between User and Caregiver | Both patients and caregivers work together using the digital app. | Highlight the important role caregivers play in the adaptation process and find ways to include their input in the adaptation procedures. | |
| G7: User Involvement in Adaptations-Assessing User Capability | Assess whether users are physically or mentally capable of handling additional burdens imposed by the system during adaptation. | Avoid overloading users and causing harm through different adjustments. | ”My parents would never consider the option of adapting the app because it is beyond their ability.” (IP 1) ”When I first start using an app, I always customise everything that can be customised. ” (FP 7) |
| (EASY) G7: Considering User’s Ability in Participating in the Adaptation | Check if users can handle additional tasks from the system during adaptation, both physically and mentally. | Ensure that users do not feel overwhelmed or harmed by introducing various changes. | |
| G8: User Involvement in Adaptations-Assessing User’s Willingness | Users exhibit different degrees of engagement in adaptation, which can be shaped by personality, cultural background, and other contextual factors. | Empower users by giving them the choice to either actively participate in shaping the adaptation or passively accept the adaptation. | ”I customise all apps, I will explore how I can adapt the app to my needs (Active involvement).” (FP 7) ”I haven’t even looked into if I can adapt the app or not (Passive involvement).” (FP 2) ”I don’t want to understand or be told what the adaptation is (Passive involvement).” (FP 1) |
| (EASY) G8: Considering User’s Willingness in Participating in the Adaptation | Some users are more involved in making changes to the app than others. This can be influenced by their personality, culture, and other factors. | Give users the option to either actively shape the adaptation or simply accept it as is. |
| Guideline | Definition | Purpose | Example |
| G9: Adjusting Granularity of the Adaptation | Degree or level of detail at which adaptations can be made to an interface. | Strike a balance in granularity to enable users to make meaningful adaptations to the user interface without overwhelming them with extensive adaptations or fatiguing them with minimal adaptations. | ”Once I understand what I am doing, I might want to push that button and get that next level of understanding(adaptation).” (IP 3) ”I think all adaptations should be implemented when users initially log into the app and change the whole app to user needs.” (FP 4,6) |
| (EASY) G9: Changing the Scale of Adaptation in the Interface | The extent to which changes can be made in an interface. | Ensure that users can make meaningful changes to the user interface without overwhelming them with too many changes or leaving them uninterested with too few changes. |
6.2. Evaluation of our Guidelines
We carried out a preliminary evaluation of our guidelines, primarily concentrating on assessing their clarity and perceived usefulness. Our objective was to determine whether target guideline users could effectively comprehend and distinguish between and make use of our guidelines. Given the possibility of limited documentation on design considerations and guidelines, particularly in domains involving new technology, it is imperative to maintain user involvement throughout the process (Hermawati and Lawson, 2015; Zaphiris et al., 2007; Lechner et al., 2013; Zaphiris et al., 2009). Although the guidelines were originally created for developers, designers, and researchers, it is crucial to ensure their organisation, usability, and significance for end users. (Lechner et al., 2013; Zaphiris et al., 2009). Therefore, we strived to further validate the efficacy and acceptance of these guidelines with the same cohort of users involved in our interview and focus group study, aiming to uncover any challenges that may not have surfaced previously and to determine whether we have accurately interpreted their perspectives, presenting them in a way that avoids misunderstandings. We reached out to all participants via email, inviting them to complete a feedback survey. Before administering the feedback survey, participants were briefed on the findings of the user study and provided with an overview of the generated guidelines. To maintain anonymity and streamline the process, participants were required to enter the unique withdrawal code provided during their initial user study registration, allowing us to correlate their responses with both the feedback survey and the user study. Subsequently, participants were requested to assess each of the guidelines and offer suggestions regarding additions, removals, or edits for each guideline. In total, 20 out of 22 participants responded to our feedback survey.
For the purpose of clarification, we designate the original versions of the guidelines as ”GO” (Guideline Original) and the updated versions as ”G”. Figure 11 shows the development of the structure of the guidelines. Initially, the framework table contained seven guidelines organised into three columns: Guidelines, Definition, and Examples. Upon analysis of the qualitative data, we integrated the purpose of the guidelines while reformulating them, leading to a revised structure with Guidelines, Definition, Purpose, and Examples. To enhance clarity, we introduced an ”EASY” version of the guideline alongside the existing guidelines, rather than removing or altering technical terms. Two guidelines (GO2 and GO5) were subdivided into more detailed ones. One guideline (GO6) was removed because it offered limited guidance and overlapped with other guidelines. These modifications resulted in a framework comprising nine guidelines, as detailed in Table 10.
6.3. Clarity of the guidelines
Although most participants found our guidelines understandable (see Figure 12), a few encountered difficulties with certain technical terms, particularly ”Granularity” (GO7) and ”Autonomy” (GO1). These terms may require further clarification or simplification to ensure broader comprehension among users. One participant in particular found it challenging to understand the guidelines, which significantly affected her ability to provide feedback. As a result, she refused to complete the feedback survey and commented:
[…]
Meanwhile, a critique regarding the difficulty in relating to the examples provided for each guideline underlines the need for clearer and more context-specific illustrations. Participants expressed confusion about how the examples applied to their own experiences and situations, particularly in the case of the Alignment of the adaptation (GO3) guideline (FP 4,7). This discrepancy in understanding could be attributed to the diverse backgrounds and perspectives of the participants. The challenge of creating relevant and accessible examples is a common issue in guideline development, as noted in the literature (Ning et al., 2019). To improve the comprehension of the example, we extracted the sample from our user study as shared by the participants themselves (see Table 10).
Rather than simply replacing terms and expressions with plain language, we acknowledge that achieving clarity and understanding across diverse user groups with varying domain knowledge may pose challenges. Research has explored how different domains can influence individuals’ perspectives on human values in eHealth applications (Obie et al., 2021). It is essential to recognise that the guidelines were crafted primarily for designers experienced in application development, whereas users may perceive and value aspects differently based on their unique perspectives and experiences. Thus, tailoring communication and guidance to align with users’ diverse viewpoints and preferences becomes imperative for effective implementation and user engagement. Therefore, we introduced an additional version of each guideline labelled as ”EASY” in Table 10. This version utilises simple language to describe the name, definition, and purpose of each guideline.
Apart from the clarity of certain guidelines, some guidelines like Alignment of the adaptation (GO2) and User involvement in the adaptation (GO5) have attracted confusion because they can be too broad and difficult to interpret (IP 1,2,6, FP 7,9). The necessity of subdividing these two guidelines into smaller and more granular components is evident. This need is further supported by Figure 12, which illustrates that GO2 and GO5 received the fewest strong agreements on their clarity. Consequently, we have divided the Alignment of the adaptation (GO2) guideline into (see Table 10):
G3: Aligning Adaptations with User’s Chronic Disease
G4: Aligning Adaptations with App Usage Patterns
G5: Aligning Adaptations with User’s Coping Style
User involvement in the adaptation (GO5) is break into (see Table 10):
G7: User Involvement in Adaptations-Assessing User Capability
G8: User Involvement in Adaptations-Assessing User’s Willingness
One participant was uncertain about the wording of the guideline, as it appeared more like a collection of themes (FP 13). The specific guidance of the Timing of the adaptation (G06) is being questioned because participants perceive it as too general, which affects its clarity. In response, we have revised the guideline by clearly framing each individual guideline. Upon further examination, it is noted that there is an overlap between Aligning Adaptations with App Usage Patterns (G4) and Timing of the adaptation (GO6). As a result, GO6 has been removed from our guidelines.
6.4. Usefulness of the guidelines
When evaluating the usefulness of the guidelines, participants expressed difficulty in relating to some of them and struggled to understand their purpose (FP 4,6, IP 3,7). They expected us to provide a rationale behind each guideline to better understand its significance. It is a common issue in guideline development that rules often specify what needs to be done without clarifying the underlying rationale or providing guidance on how to implement them effectively (Carter, 1999). As a response, we have supplemented each guideline with its purpose to offer context and aid in understanding. Participants with experience in application design noted the potential overlap between our proposed guidelines and existing general guidelines for designing mHealth solutions (FP 10,13,14). FP14 noted that:
[…]
This observation of overlap in the guidelines resonates with findings from other research studies, which have also identified this as a common issue (Srivastava et al., 2021). We refined the guidelines by providing clear purposes for each guideline, our objective is to help researchers and practitioners better understand the challenges faced by users with chronic diseases and how each guideline seeks to mitigate them. Participants in the evaluation study emphasised the need to prioritise certain guidelines within our framework based on the chronic disease domain. FP8 highlighted the importance of ”certain guidelines over others”, while IP5 noted that ”some guidelines seemed less relevant to their needs”. This emphasises the challenge of designing for the hypothetical ”general” user (Carter, 1999). However, participants in our feedback survey have used different types of applications related to the domain of chronic diseases. For instance, younger users may use specific apps to monitor their health measurements, with minimal involvement from their caregivers, making the guidelines for user-caregiver collaboration less effective. Additionally, the achievement of all design goals for a computer-based product or service often involves trade-offs (Johnson, 2020). We have discussed the trade-offs associated with the sets of strategies identified in Section 4.5. The inherent nature of trade-offs is inevitable in our guidelines. For example, consider the trade-off between privacy and personalisation in the design of an mHealth application. While robust privacy measures are necessary to protect sensitive health information, excessive restrictions on data collection and sharing may limit the app’s ability to adapt the UI to suit individual users’ health needs. In such cases, designers must carefully consider the implications of each decision and prioritise the guidelines that best align with the app’s goals and target user demographics (Ning et al., 2019). When faced with conflicting rules or multiple guidelines, designers must weigh their significance and make informed decisions to effectively address design challenges.
7. Threats to Validity
External validity. For the interview and focus group study, we employed STGT’s theoretical sampling technique, which is aimed at generating insights from representative samples rather than achieving generalisability (Hargittai and Micheli, 2019; Hoda, 2021). Therefore, the qualitative findings of our study cannot be generalised to broader populations, but offer valuable insights into how users perceive AUIs in mHealth applications for individuals with chronic diseases. Future research endeavours should seek to validate our findings with more diverse populations, particularly those with lower levels of digital literacy (Hargittai and Micheli, 2019). Secondly, the use of a high-fidelity design prototype in our user study may have inadvertently introduced discrepancies between the prototyped adaptations and the practical applications. This prototype may not fully capture the intricacies of real-world interactions, such as the dynamic interaction between adaptations and other system functionalities or fail to account for users’ nuanced allocation of time to specific adaptations (Bunt et al., 2004; Gajos et al., 2006, 2008). These complexities are crucial to consider as they could influence users’ perceptions and behaviours in ways that the prototype might not accurately reflect. Therefore, future research efforts should explore ways to bridge this gap between high-fidelity prototypes and real-world implementations to ensure that the findings are applicable and actionable in practical settings. Thirdly, we observed demographic differences in our study concerning nationality, age, and the clinic-based population. Although we shared the survey on social media to obtain worldwide participation, we could not achieve it and found that the majority of the participants (49%) were from Australia (Table 4). Hence, similar to the study (Robbins et al., 2017), our findings may not be generalisable to the global population of individuals with chronic diseases. Our survey sample exhibited a skew towards younger demographics compared to the participants involved in the interview and focus group studies. This demographic disparity is significant as patterns of usage and preference are likely to vary across different age groups, with older populations potentially exhibiting different behaviours and preferences. Additionally, it is important to acknowledge that our study surveyed a general population rather than specifically targeting individuals known to be in hospital or clinic-based settings. These individuals, who regularly engage with healthcare services, may exhibit different behaviours and usage patterns of mHealth applications compared to those of the general population with milder symptoms.
Internal validity. Within our prototype, we curated a diverse range of adaptations. However, participants may have encountered instances of confusion when interacting with certain elements of the AUIs. To mitigate this issue, we incorporated an explanatory video to aid participants in navigating through the various adaptations. Furthermore, we supplemented this with detailed instructions for reference, ensuring participants had guidance readily available should they encounter any challenges while using the system. Similarly, we anticipated potential complexities during the survey phase, and thus provided two concrete examples of AUIs to clarify the AUI concepts for participants. We implemented the STGT approach for qualitative data analysis and facilitated extensive team discussions to review and refine our analyses, findings, and presentation, thus mitigating potential biases. Although offering compensation to participants for their participation in interviews and focus group studies can raise concerns about the potential of participants to provide false information to qualify for the study (Head, 2009), it is noteworthy that four participants declined compensation. Instead, they expressed a preference for the funds to be allocated towards further research endeavours. This underscores the voluntary nature of our study and emphasises the perceived importance of the research from the perspective of participants. The survey responses regarding opinions on adaptations for mHealth apps may shift based on their evolving needs and experiences as users interact with mHealth apps over time, and they may also stop using the app, as long-term use of the mHealth app has always been an issue (Woldaregay et al., 2018).
Conclusion validity. While our study focuses on various chronic diseases, it is important to note that the findings may have limitations in terms of generalisability. Nevertheless, our approach aligns with the common practice of including individuals with multiple chronic diseases in research (Elzen et al., 2007). In fact, some participants in our study presented with multiple chronic diseases, mirroring real-life scenarios where individuals often contend with comorbidity alongside their primary condition (di San Marco et al., 2019; Islam et al., 2014). Moving forward, the practical utility of our guidelines warrants validation through user evaluations of adaptive mHealth apps. Our study represents a preliminary evaluation of the framework, and a larger-scale follow-up study involving industry practitioners could provide valuable insights. It is essential to recognise that our framework is evolutionary in nature, and we encourage further research to explore its applicability to population groups at the intersections and with unique needs.
8. conclusion
In response to society’s demand for improved mHealth applications to assist diverse users in managing chronic diseases, our study explored user perspectives on various Adaptive User Interface (AUI) approaches in chronic disease management apps. Through this research, we identified significant challenges associated with the use of AUIs and proposed strategies to enhance adaptation design through interviews and focus group studies. Our analysis was further supported by a post-session survey, which corroborated the emergence of these categories. However, it is essential to acknowledge that the adoption of these strategies may entail certain costs, influenced by variables such as user involvement levels, prior experience with mHealth apps, and individual health conditions. Furthermore, we conducted an in-depth exploration of the relationship between specific challenges and the corresponding strategies proposed, illuminating the nuanced trade-offs involved. This provided valuable insights into prioritising user and app characteristics for optimal adaptation design. The quantitative survey provided additional key information on emerging categories and revealed participants’ preferences for different adaptations, data types, collection methods, and levels of engagement. Finally, these insights and categories were synthesised and mapped to existing guidelines or design considerations for mHealth app design adaptation to distil a set of guidelines. We refined these guidelines with 20 participants. These guidelines, which serve as a roadmap for the design of future AUIs in mHealth apps targeting chronic disease development, align with existing design principles, ensuring their relevance and applicability in real-world settings. We characterise this effort in refining guidelines as a preliminary evaluation, and a comprehensive follow-up investigation involving professionals from the industry and academic could be the subsequent phase. It is important to note that our framework is progressive and we advocate for additional research to expand or examine its applicability to different demographic groups.
Acknowledgements.
Wang, Grundy and Madugalla are supported by ARC Laureate Fellowship FL190100035. We would like to acknowledge Daniel Gaspar Figueiredo, Elton Lobo, Michael Wheeler, and Paul Jansons for helping us recruit participants for the user study. Special thanks to all the participants for sharing their valuable perspectives and experiences. The following colleagues provided valuable assistance in the form of comments on earlier drafts, data analysis, prototype evaluation and/or graphics in the study: Dulaji Hidellaarachchi, Elizabeth Manias, Kashumi Madampe and Weimin Wang. We are grateful to Rashina Hoad for her assistance in applying Socio-Technical Grounded Theory (STGT) for data analysis in our paper.References
- (1)
- Abrahão et al. (2021) Silvia Abrahão, Emilio Insfran, Arthur Sluÿters, and Jean Vanderdonckt. 2021. Model-based intelligent user interface adaptation: challenges and future directions. Software and Systems Modeling 20, 5 (2021), 1335–1349.
- Acocella (2012) Ivana Acocella. 2012. The focus groups in social research: advantages and disadvantages. Quality & Quantity 46 (2012), 1125–1136.
- Akbar et al. (2020) Saba Akbar, Enrico Coiera, and Farah Magrabi. 2020. Safety concerns with consumer-facing mobile health applications and their consequences: a scoping review. Journal of the American Medical Informatics Association 27, 2 (2020), 330–340.
- Akiki et al. (2014) Pierre A. Akiki, Arosha K. Bandara, and Yijun Yu. 2014. Adaptive Model-Driven User Interface Development Systems. ACM Computing Surveys (CSUR) 47, 1 (may 2014). https://doi.org/10.1145/2597999
- Albahri et al. (2018) Ahmed Shihab Albahri, AA Zaidan, Osamah Shihab Albahri, BB Zaidan, and MA Alsalem. 2018. Real-time fault-tolerant mHealth system: Comprehensive review of healthcare services, opens issues, challenges and methodological aspects. Journal of medical systems 42 (2018), 1–56.
- Albers (1997) Michael J Albers. 1997. Cognitive strain as a factor in effective document design. In Proceedings of the 15th annual international conference on Computer documentation. 1–6.
- Aljedaani et al. (2021) Bakheet Aljedaani, M Ali Babar, et al. 2021. Challenges with developing secure mobile health applications: Systematic review. JMIR mHealth and uHealth 9, 6 (2021), e15654.
- Alvarez-Cortes et al. (2007) Victor Alvarez-Cortes, Benjamin E Zayas-Perez, Victor Huga Zarate-Silva, and Jorge A Ramirez Uresti. 2007. Current trends in adaptive user interfaces: Challenges and applications. In Electronics, Robotics and Automotive Mechanics Conference (CERMA 2007). IEEE, 312–317.
- Anderson et al. (2016) Kevin Anderson, Oksana Burford, and Lynne Emmerton. 2016. Mobile health apps to facilitate self-care: a qualitative study of user experiences. PloS one 11, 5 (2016), e0156164.
- Attfield et al. (2011) S Attfield, G Kazai, M Lalmas, and B Piwowarski. 2011. Towards a science of user engagement. WSDM workshop on user modelling for web applications.
- Audulv (2013) Åsa Audulv. 2013. The over time development of chronic illness self-management patterns: a longitudinal qualitative study. BMC public health 13, 1 (2013), 1–15.
- Awada et al. (2018) Imad Alex Awada, Irina Mocanu, Dumitru-Iulian Nastac, Dan Benta, and Serban Radu. 2018. Adaptive user interface for healthcare application for people with dementia. In 2018 17th RoEduNet Conference: Networking in Education and Research (RoEduNet). IEEE, 1–5.
- Baroni et al. (2022) Ilaria Baroni, Gloria Re Calegari, Damiano Scandolari, and Irene Celino. 2022. AI-TAM: a model to investigate user acceptance and collaborative intention in human-in-the-loop AI applications. Human Computation 9, 1 (2022), 1–21.
- Beaglehole et al. (2008) Robert Beaglehole, JoAnne Epping-Jordan, Vikram Patel, Mickey Chopra, Shah Ebrahim, Michael Kidd, and Andy Haines. 2008. Improving the prevention and management of chronic disease in low-income and middle-income countries: a priority for primary health care. The Lancet 372, 9642 (2008), 940–949.
- Bentley et al. (2014) Claire L Bentley, Lauren A Powell, Alison Orrell, and Gail A Mountain. 2014. Addressing design and suitability barriers to telecare use: has anything changed? Technology and Disability 26, 4 (2014), 221–235.
- Bettman (1971) James R Bettman. 1971. The structure of consumer choice processes. Journal of marketing research 8, 4 (1971), 465–471.
- Betts et al. (2019) Lucy R Betts, Rowena Hill, and Sarah E Gardner. 2019. “There’s not enough knowledge out there”: Examining older adults’ perceptions of digital technology use and digital inclusion classes. Journal of Applied Gerontology 38, 8 (2019), 1147–1166.
- Billsus and Pazzani (2007) Daniel Billsus and Michael J Pazzani. 2007. Adaptive news access. In The adaptive web: methods and strategies of web personalization. Springer, 550–570.
- Bricca et al. (2022) Alessio Bricca, Alessandro Pellegrini, Graziella Zangger, Jonas Ahler, Madalina Jäger, Søren T Skou, et al. 2022. The quality of health apps and their potential to promote behavior change in patients with a chronic condition or multimorbidity: systematic search in app store and Google play. JMIR mHealth and uHealth 10, 2 (2022), e33168.
- Browne et al. (1990) Dermot Browne, Mike Norman, and Dave Riches. 1990. Why build adaptive systems? In Adaptive user interfaces. Elsevier, 15–57.
- Bunt et al. (2004) Andrea Bunt, Cristina Conati, and Joanna McGrenere. 2004. What role can adaptive support play in an adaptable system?. In IUI2004. 117–124.
- Bunt et al. (2012) Andrea Bunt, Matthew Lount, and Catherine Lauzon. 2012. Are explanations always important? A study of deployed, low-cost intelligent interactive systems. In Proceedings of the 2012 ACM international conference on Intelligent User Interfaces. 169–178.
- Burke et al. (2009) James William Burke, MDJ McNeill, Darryl K Charles, Philip J Morrow, Jacqui H Crosbie, and Suzanne M McDonough. 2009. Optimising engagement for stroke rehabilitation using serious games. The Visual Computer 25 (2009), 1085–1099.
- Cajita et al. (2018) Maan Isabella Cajita, Nancy A Hodgson, Katherine Wai Lam, Sera Yoo, and Hae-Ra Han. 2018. Facilitators of and barriers to mHealth adoption in older adults with heart failure. Computers, informatics, nursing: CIN 36, 8 (2018), 376.
- Camacho-Rivera et al. (2020) Marlene Camacho-Rivera, Jessica Yasmine Islam, Argelis Rivera, Denise Christina Vidot, et al. 2020. Attitudes toward using COVID-19 mHealth tools among adults with chronic health conditions: secondary data analysis of the COVID-19 impact survey. JMIR mHealth and uHealth 8, 12 (2020), e24693.
- Carter (1999) Jim Carter. 1999. Incorporating standards and guidelines in an approach that balances usability concerns for developers and end users. Interacting with Computers 12, 2 (1999), 179–206.
- Cassileth et al. (1980) Barrie R. Cassileth, Robert V. Zupkis, K Sutton-Smith, and Vicki March. 1980. Information and participation preferences among cancer patients. Annals of internal medicine 92 6 (1980), 832–6. https://api.semanticscholar.org/CorpusID:9220754
- Cheverst et al. (2005) Keith Cheverst, Hee Eon Byun, Dan Fitton, Corina Sas, Chris Kray, and Nicolas Villar. 2005. Exploring issues of user model transparency and proactive behaviour in an office environment control system. User Modeling and User-Adapted Interaction 15 (2005), 235–273.
- Chow et al. (2016) Clara K Chow, Nilshan Ariyarathna, Sheikh Mohammed Shariful Islam, Aravinda Thiagalingam, and Julie Redfern. 2016. mHealth in cardiovascular health care. Heart, Lung and Circulation 25, 8 (2016), 802–807.
- Codish and Ravid (2014) David Codish and Gilad Ravid. 2014. Personality based gamification-Educational gamification for extroverts and introverts. In Proceedings of the 9th CHAIS Conference for the Study of Innovation and Learning Technologies: Learning in the Technological Era, Vol. 1. The Open University of Israel Ra’anana, 36–44.
- Cristea and Calvi (2003) Alexandra Cristea and Licia Calvi. 2003. The three layers of adaptation granularity. In International Conference on User Modeling. Springer, 4–14.
- Cruz-Ramos et al. (2022) Nancy Aracely Cruz-Ramos, Giner Alor-Hernández, Luis Omar Colombo-Mendoza, José Luis Sánchez-Cervantes, Lisbeth Rodríguez-Mazahua, and Luis Rolando Guarneros-Nolasco. 2022. mHealth apps for self-management of cardiovascular diseases: a scoping review. In Healthcare, Vol. 10. MDPI, 322.
- Csikszentmihalyi et al. (2014) Mihaly Csikszentmihalyi, Martin EP Seligman, and Mihaly Csikszentmihalyi. 2014. Positive psychology: An introduction. Flow and the Foundations of Positive Psychology: The Collected Works of Mihaly Csikszentmihalyi (2014), 279–298.
- Deci and Ryan (2013) Edward L Deci and Richard M Ryan. 2013. Intrinsic motivation and self-determination in human behavior. Springer Science & Business Media.
- Deiss et al. (2006) Dorothee Deiss, Jan Bolinder, Jean-Pierre Riveline, Tadej Battelino, Emanuele Bosi, Nadia Tubiana-Rufi, David Kerr, and Moshe Phillip. 2006. Improved glycemic control in poorly controlled patients with type 1 diabetes using real-time continuous glucose monitoring. Diabetes care 29, 12 (2006), 2730–2732.
- Denzin (2017) Norman K Denzin. 2017. The research act: A theoretical introduction to sociological methods. Transaction publishers.
- Dessart et al. (2011) Charles-Eric Dessart, Vivian Genaro Motti, and Jean Vanderdonckt. 2011. Showing user interface adaptivity by animated transitions. In Proceedings of the 3rd ACM SIGCHI symposium on Engineering interactive computing systems. 95–104.
- Deuschel (2018) Tilman Deuschel. 2018. On the Influence of Human Factors in Adaptive User Interface Design. In Adjunct Publication of the 26th Conference on User Modeling, Adaptation and Personalization. 187–190.
- Deuschel and Scully (2016) Tilman Deuschel and Ted Scully. 2016. On the Importance of Spatial Perception for the Design of Adaptive User Interfaces. 2016 IEEE 10th International Conference on Self-Adaptive and Self-Organizing Systems (SASO) (2016), 70–79. https://api.semanticscholar.org/CorpusID:8789842
- di San Marco et al. (2019) Eugenia Cao di San Marco, Elena Vegni, and Lidia Borghi. 2019. Chronic Illnesses, Vulnerability, and Uncertainty: How Do Recent Challenges Impact on Patient-Centered Medicine? International Journal of Patient-Centered Healthcare (IJPCH) 9, 1 (2019), 50–63.
- Duplaga (2015) Mariusz Duplaga. 2015. A cross-sectional study assessing determinants of the attitude to the introduction of eHealth services among patients suffering from chronic conditions. BMC medical informatics and decision making 15, 1 (2015), 1–15.
- Eiband et al. (2021) Malin Eiband, Daniel Buschek, and Heinrich Hussmann. 2021. How to support users in understanding intelligent systems? Structuring the discussion. In 26th International Conference on Intelligent User Interfaces. 120–132.
- Elzen et al. (2007) Henrike Elzen, Joris PJ Slaets, Tom AB Snijders, and Nardi Steverink. 2007. Evaluation of the chronic disease self-management program (CDSMP) among chronically ill older people in the Netherlands. Social science & medicine 64, 9 (2007), 1832–1841.
- Eslami et al. (2018) Mahboubeh Eslami, Mohammad Firoozabadi, and Elaheh Homayounvala. 2018. User preferences for adaptive user interfaces in health information systems. Universal Access in the Information Society 17 (2018), 875–883.
- Evers and Day (1997) Vanessa Evers and Donald Day. 1997. The role of culture in interface acceptance. In Human-Computer Interaction INTERACT’97: IFIP TC13 International Conference on Human-Computer Interaction, 14th–18th July 1997, Sydney, Australia. Springer, 260–267.
- Fazio and Zanna (1978) Russell H Fazio and Mark P Zanna. 1978. Attitudinal qualities relating to the strength of the attitude-behavior relationship. Journal of experimental social psychology 14, 4 (1978), 398–408.
- Findlater and McGrenere (2004) Leah Findlater and Joanna McGrenere. 2004. A comparison of static, adaptive, and adaptable menus. CHI2004 (2004). https://api.semanticscholar.org/CorpusID:207742233
- Findlater and McGrenere (2007) Leah Findlater and Joanna McGrenere. 2007. Evaluating reduced-functionality interfaces according to feature findability and awareness. In IFIP Conference on Human-Computer Interaction. Springer, 592–605.
- Findlater and McGrenere (2010) Leah Findlater and Joanna McGrenere. 2010. Beyond performance: Feature awareness in personalized interfaces. Int. J. Hum. Comput. Stud. 68 (2010), 121–137. https://api.semanticscholar.org/CorpusID:10129897
- Fischer (2006) Gerhard Fischer. 2006. Distributed intelligence: extending the power of the unaided, individual human mind. In Proceedings of the working conference on Advanced visual interfaces. 7–14.
- Folkman and Lazarus (1980) Susan Folkman and Richard S Lazarus. 1980. An analysis of coping in a middle-aged community sample. Journal of health and social behavior (1980), 219–239.
- Fröhlich et al. (2009) Nadine Fröhlich, Andreas Meier, Thorsten Möller, Marco Savini, Heiko Schuldt, and Joël Vogt. 2009. LoCa–towards a context-aware infrastructure for eHealth applications. In Proc. of the 15th Int’l Conference on Distributed Multimedia Systems (DMS’09). Citeseer.
- Fusch et al. (2018) Patricia Fusch, Gene E Fusch, and Lawrence R Ness. 2018. Denzin’s paradigm shift: Revisiting triangulation in qualitative research. Journal of Sustainable Social Change 10, 1 (2018), 2.
- Gajos and Chauncey (2017) Krzysztof Z Gajos and Krysta Chauncey. 2017. The Influence of Personality Traits and Cognitive Load on the Use of Adaptive User Interfaces. Proceedings of the 22nd International Conference on Intelligent User Interfaces (2017). https://api.semanticscholar.org/CorpusID:10739831
- Gajos et al. (2006) Krzysztof Z Gajos, Mary Czerwinski, Desney S Tan, and Daniel S Weld. 2006. Exploring the design space for adaptive graphical user interfaces. In AVI2006. 201–208.
- Gajos et al. (2008) Krzysztof Z Gajos, Katherine Everitt, Desney S Tan, Mary Czerwinski, and Daniel S Weld. 2008. Predictability and accuracy in adaptive user interfaces. In CHI2008. 1271–1274.
- Glaser (1992) Barney G Glaser. 1992. Basics of grounded theory analysis: Emergence vs forcing. (No Title) (1992).
- Glass et al. (2008) Alyssa Glass, Deborah L McGuinness, and Michael Wolverton. 2008. Toward establishing trust in adaptive agents. In IUI2008. 227–236.
- Gómez-Carmona et al. (2024) Oihane Gómez-Carmona, Diego Casado-Mansilla, Diego López-de Ipiña, and Javier García-Zubia. 2024. Human-in-the-loop machine learning: Reconceptualizing the role of the user in interactive approaches. Internet of Things (2024), 101048.
- Greenwood et al. (2003) Sue Greenwood, John Nealon, and Peter Marshall. 2003. Agent-based user Interface Adaptivity in a medical decision support system. Applications of Software Agent Technology in the Health Care Domain (2003), 35–47.
- Grua et al. (2020) Eoin Martino Grua, Martina De Sanctis, and Patricia Lago. 2020. A reference architecture for personalized and self-adaptive e-health apps. In European Conference on Software Architecture. Springer, 195–209.
- Gullà et al. (2019) Francesca Gullà, Roberto Menghi, and Michele Germani. 2019. Study of the Usability of an Adaptive Smart Home Interface for People with Alzheimer’s Disease. In Ambient Assisted Living: Italian Forum 2017 8. Springer, 261–269.
- Haggag et al. (2022) Omar Haggag, John Grundy, Mohamed Abdelrazek, and Sherif Haggag. 2022. A large scale analysis of mHealth app user reviews. Empirical Software Engineering 27, 7 (2022), 196.
- Hallberg et al. (2016) Inger Hallberg, Agneta Ranerup, and K Kjellgren. 2016. Supporting the self-management of hypertension: Patients’ experiences of using a mobile phone-based system. Journal of human hypertension 30, 2 (2016), 141–146.
- Hamine et al. (2015) Saee Hamine, Emily Gerth-Guyette, Dunia Faulx, Beverly B Green, and Amy Sarah Ginsburg. 2015. Impact of mHealth chronic disease management on treatment adherence and patient outcomes: a systematic review. Journal of medical Internet research 17, 2 (2015), e52.
- Hammond (1987) Nick Hammond. 1987. Principles from the psychology of skill acquisition. In Applying cognitive psychology to user-interface design. 163–187.
- Han et al. (2010) Jung Hoon Han, Naomi Sunderland, Elizabeth Kendall, Ori Gudes, and Garth Henniker. 2010. Professional practice and innovation: Chronic disease, geographic location and socioeconomic disadvantage as obstacles to equitable access to e-health. Health Information Management Journal 39, 2 (2010), 30–36.
- Hargittai and Micheli (2019) Eszter Hargittai and Marina Micheli. 2019. Internet skills and why they matter. Society and the internet: How networks of information and communication are changing our lives 109 (2019), 109–124.
- Harman et al. (2014) Mark Harman, Yue Jia, William B Langdon, Justyna Petke, Iman Hemati Moghadam, Shin Yoo, and Fan Wu. 2014. Genetic improvement for adaptive software engineering (keynote). In 9th International Symposium on Software Engineering for Adaptive and Self-Managing Systems. 1–4.
- Harper et al. (2015) F Maxwell Harper, Funing Xu, Harmanpreet Kaur, Kyle Condiff, Shuo Chang, and Loren Terveen. 2015. Putting users in control of their recommendations. In 9th ACM Conference on Recommender Systems. 3–10.
- Harris (2021) Jenine K Harris. 2021. Primer on binary logistic regression. Family Medicine and Community Health 9, Suppl 1 (2021).
- Harris and Brown (2019) Lois R Harris and Gavin TL Brown. 2019. Mixing interview and questionnaire methods: Practical problems in aligning data. Practical Assessment, Research, and Evaluation 15, 1 (2019), 1.
- Harvey et al. (2012) Alison Harvey, Angela Brand, Stephen T Holgate, Lars V Kristiansen, Hans Lehrach, Aarno Palotie, and Barbara Prainsack. 2012. The future of technologies for personalised medicine. New biotechnology 29, 6 (2012), 625–633.
- Head (2009) Emma Head. 2009. The ethics and implications of paying participants in qualitative research. International Journal of Social Research Methodology 12, 4 (2009), 335–344.
- Hermawati and Lawson (2015) Setia Hermawati and Glyn Lawson. 2015. A user-centric methodology to establish usability heuristics for specific domains. Contemp. Ergon. Hum. Factors 2015 (2015), 80–85.
- Hird et al. (2016) Nick Hird, Samik Ghosh, and Hiroaki Kitano. 2016. Digital health revolution: perfect storm or perfect opportunity for pharmaceutical R&D? Drug discovery today 21, 6 (2016), 900–911.
- Hoda (2021) Rashina Hoda. 2021. Socio-technical grounded theory for software engineering. IEEE Transactions on Software Engineering 48, 10 (2021), 3808–3832.
- Höök (1998) Kristina Höök. 1998. Tutorial 2: Designing and evaluating intelligent user interfaces. In IUI1998. 5–6.
- Höök (2000) Kristina Höök. 2000. Steps to take before intelligent user interfaces become real. Interacting with computers 12, 4 (2000), 409–426.
- Inbar and Tractinsky (2009) Ohad Inbar and Noam Tractinsky. 2009. FEATURE The incidental user. Interactions 16, 4 (2009), 56–59.
- Islam et al. (2014) M Mofizul Islam, Jose M Valderas, Laurann Yen, Paresh Dawda, Tanisha Jowsey, and Ian S McRae. 2014. Multimorbidity and comorbidity of chronic diseases among the senior Australians: prevalence and patterns. PloS one 9, 1 (2014), e83783.
- Jabeen et al. (2019) Farzana Jabeen, Linmi Tao, Yirou Guo, Shiyu Zhang, and Shanshan Mei. 2019. Improving Mobile Device interaction for Parkinson’s Disease Patients via PD-Helper.. In SEKE. 529–692.
- Jameson (2007) Anthony Jameson. 2007. Adaptive interfaces and agents. In The human-computer interaction handbook. CRC press, 459–484.
- Jameson and Schwarzkopf (2002) Anthony Jameson and Eric Schwarzkopf. 2002. Pros and cons of controllability: An empirical study. In International Conference on Adaptive Hypermedia and Adaptive Web-Based Systems. Springer, 193–202.
- Johnson (2020) Jeff Johnson. 2020. Designing with the mind in mind: simple guide to understanding user interface design guidelines. Morgan Kaufmann.
- Kay (2001) Judy Kay. 2001. Learner control. User modeling and user-adapted interaction 11 (2001), 111–127.
- Kayyali et al. (2017) Reem Kayyali, Aliki Peletidi, Muhammad Ismail, Zahra Hashim, Pedro Bandeira, and Jennifer Bonnah. 2017. Awareness and use of mHealth apps: a study from England. Pharmacy 5, 2 (2017), 33.
- Kc et al. (2021) Bhuvan Kc, Alian A Alrasheedy, Bey Hing Goh, Ali Blebil, Nida Sajid Ali Bangash, Mohamed Izham Mohamed Ibrahim, and Inayat Ur Rehman. 2021. The types and pattern of use of mobile health applications among the general population: A cross-sectional study from Selangor, Malaysia. Patient preference and adherence (2021), 1755–1762.
- Kelders et al. (2012) Saskia M Kelders, Robin N Kok, Hans C Ossebaard, and Julia EWC Van Gemert-Pijnen. 2012. Persuasive system design does matter: a systematic review of adherence to web-based interventions. Journal of medical Internet research 14, 6 (2012), e152.
- Kornfield et al. (2020) Rachel Kornfield, Renwen Zhang, Jennifer Nicholas, Stephen M Schueller, Scott A Cambo, David C Mohr, and Madhu Reddy. 2020. ” Energy is a Finite Resource”: Designing Technology to Support Individuals across Fluctuating Symptoms of Depression. In CHI2020. 1–17.
- Kühme (1993) Thomas Kühme. 1993. A user-centered approach to adaptive interfaces. In Proceedings of the 1st international conference on Intelligent user interfaces. 243–245.
- Kulesza et al. (2013) Todd Kulesza, Simone Stumpf, Margaret Burnett, Sherry Yang, Irwin Kwan, and Weng-Keen Wong. 2013. In 2013 IEEE Symposium on visual languages and human centric computing. IEEE, 3–10.
- Kurtz (2005) Matthew M Kurtz. 2005. Neurocognitive impairment across the lifespan in schizophrenia: an update. Schizophrenia research 74, 1 (2005), 15–26.
- Lage et al. (2014) Ricardo Lage, Peter Dolog, and Martin Leginus. 2014. The role of adaptive elements in web-based surveillance system user interfaces. In User Modeling, Adaptation, and Personalization: 22nd International Conference, UMAP 2014, Aalborg, Denmark, July 7-11, 2014. Proceedings 22. Springer, 350–362.
- Lanier (1995) Jaron Lanier. 1995. Agents of alienation. interactions 2, 3 (1995), 66–72.
- Lavie and Meyer (2010) Talia Lavie and Joachim Meyer. 2010. Benefits and costs of adaptive user interfaces. International Journal of Human-Computer Studies 68, 8 (2010), 508–524.
- Lechner et al. (2013) Bettina Lechner, Ann L Fruhling, Stacie Petter, and Harvey P Siy. 2013. The Chicken and the Pig: User Involvement in Developing Usability Heuristics.. In AMCIS.
- Lewis and Wyatt (2014) Thomas Lorchan Lewis and Jeremy C Wyatt. 2014. mHealth and mobile medical apps: a framework to assess risk and promote safer use. Journal of medical Internet research 16, 9 (2014), e3133.
- Li et al. (2020) Jiaoyang Li, Cheng Zhang, Xixi Li, and Chenghong Zhang. 2020. Patients’ emotional bonding with MHealth apps: An attachment perspective on patients’ use of MHealth applications. International Journal of Information Management 51 (2020), 102054.
- Li and Gajos (2014) Louis Li and Krzysztof Z Gajos. 2014. Adaptive click-and-cross: adapting to both abilities and task improves performance of users with impaired dexterity. IUI2014 (2014). https://api.semanticscholar.org/CorpusID:3823807
- Lin et al. (2018) Zhiyuan Lin, Tim Althoff, and Jure Leskovec. 2018. I’ll be back: on the multiple lives of users of a mobile activity tracking application. In Proceedings of the 2018 World Wide Web Conference. 1501–1511.
- Lohr (2015) Steve Lohr. 2015. The healing power of your own medical records. The New York Times 31, 3 (2015), 1–8.
- Lorig and Holman (2003) Kate R Lorig and Halsted R Holman. 2003. Self-management education: history, definition, outcomes, and mechanisms. Annals of behavioral medicine 26, 1 (2003), 1–7.
- Martin-Hammond et al. (2019) Aqueasha Martin-Hammond, Sravani Vemireddy, Kartik Rao, et al. 2019. Exploring older adults’ beliefs about the use of intelligent assistants for consumer health information management: A participatory design study. JMIR aging 2, 2 (2019), e15381.
- McDougall et al. (2000) Siné JP McDougall, Oscar De Bruijn, and Martin B Curry. 2000. Exploring the effects of icon characteristics on user performance: The role of icon concreteness, complexity, and distinctiveness. Journal of Experimental Psychology: Applied 6, 4 (2000), 291.
- McLean et al. (2011) Susannah McLean, Denis Protti, and Aziz Sheikh. 2011. Telehealthcare for long term conditions. Bmj 342 (2011).
- McTear (1993) Michael F McTear. 1993. User modelling for adaptive computer systems: a survey of recent developments. Artificial intelligence review 7, 3 (1993), 157–184.
- Melenhorst et al. (2001) Anne-Sophie Melenhorst, Wendy A Rogers, and Evan C Caylor. 2001. The use of communication technologies by older adults: exploring the benefits from the user’s perspective. In Proceedings of the human factors and ergonomics society annual meeting, Vol. 45. SAGE Publications Sage CA: Los Angeles, CA, 221–225.
- Meyer et al. (2016) Jochen Meyer, Jochen Schnauber, Wilko Heuten, Harm Wienbergen, Rainer Hambrecht, Hans-Jürgen Appelrath, and Susanne Boll. 2016. Exploring longitudinal use of activity trackers. In 2016 IEEE International Conference on Healthcare Informatics (ICHI). IEEE, 198–206.
- Mezhoudi et al. (2015) Nesrine Mezhoudi, Iyad Khaddam, and Jean Vanderdonckt. 2015. Toward usable intelligent user interface. In Human-Computer Interaction: Interaction Technologies: 17th International Conference, HCI International 2015, Los Angeles, CA, USA, August 2-7, 2015, Proceedings, Part II 17. Springer, 459–471.
- Mohan et al. (2008) P Mohan, D Marin, S Sultan, and A Deen. 2008. MediNet: personalizing the self-care process for patients with diabetes and cardiovascular disease using mobile telephony. In 2008 30th Annual International Conference of the IEEE Engineering in Medicine and Biology Society. IEEE, 755–758.
- Montague and Xu (2012) Enid Montague and Jie Xu. 2012. Understanding active and passive users: The effects of an active user using normal, hard and unreliable technologies on user assessment of trust in technology and co-user. Applied ergonomics 43, 4 (2012), 702–712.
- Mukhiya et al. (2020) Suresh Kumar Mukhiya, Jo Dugstad Wake, Yavuz Inal, and Yngve Lamo. 2020. Adaptive systems for internet-delivered psychological treatments. IEEE Access 8 (2020), 112220–112236.
- Munteanu and Salah (2017) Cosmin Munteanu and Albert Ali Salah. 2017. Multimodal technologies for seniors: challenges and opportunities. The Handbook of Multimodal-Multisensor Interfaces: Foundations, User Modeling, and Common Modality Combinations-Volume 1 (2017), 319–362.
- Nemes et al. (2009) Szilard Nemes, Junmei Miao Jonasson, Anna Genell, and Gunnar Steineck. 2009. Bias in odds ratios by logistic regression modelling and sample size. BMC medical research methodology 9 (2009), 1–5.
- Ning et al. (2019) Weining Ning, Joy Goodman-Deane, and P John Clarkson. 2019. Addressing cognitive challenges in design–a review on existing approaches. In Proceedings of the Design Society: International Conference on Engineering Design, Vol. 1. Cambridge University Press, 2775–2784.
- Norcio and Stanley (1989) Anthony F. Norcio and Jaki Stanley. 1989. Adaptive Human-Computer Interfaces: A Literature Survey and Perspective. IEEE Transactions on Systems, Man and Cybernetics 19, 2 (1989), 399–408. https://doi.org/10.1109/21.31042
- Obie et al. (2021) Humphrey O Obie, Mojtaba Shahin, John Grundy, Burak Turhan, Li Li, Waqar Hussain, and Jon Whittle. 2021. Does domain change the opinion of individuals on human values? A preliminary investigation on eHealth apps end-users. In 2021 28th Asia-Pacific Software Engineering Conference (APSEC). IEEE, 531–535.
- O’Brien and Toms (2008) Heather L O’Brien and Elaine G Toms. 2008. What is user engagement? A conceptual framework for defining user engagement with technology. Journal of the American society for Information Science and Technology 59, 6 (2008), 938–955.
- Oertel et al. (2020) Catharine Oertel, Ginevra Castellano, Mohamed Chetouani, Jauwairia Nasir, Mohammad Obaid, Catherine Pelachaud, and Christopher Peters. 2020. Engagement in human-agent interaction: An overview. Frontiers in Robotics and AI 7 (2020), 92.
- Oppenheim (2000) Abraham Naftali Oppenheim. 2000. Questionnaire design, interviewing and attitude measurement. Bloomsbury Publishing.
- Pagiatakis et al. (2020) Catherine Pagiatakis, David Rivest-Hénault, David Roy, Francis Thibault, and Di Jiang. 2020. Intelligent interaction interface for medical emergencies: Application to mobile hypoglycemia management. Smart Health 15 (2020), 100091.
- Parameswaran et al. (2015) Srikanth Parameswaran, Rajiv Kishore, and Pu Li. 2015. Within-study measurement invariance of the UTAUT instrument: An assessment with user technology engagement variables. Information & management 52, 3 (2015), 317–336.
- Paré et al. (2007) Guy Paré, Mirou Jaana, and Claude Sicotte. 2007. Systematic review of home telemonitoring for chronic diseases: the evidence base. Journal of the American Medical Informatics Association 14, 3 (2007), 269–277.
- Parker et al. (2019) Charlie Parker, Sam Scott, and Alistair Geddes. 2019. Snowball sampling. SAGE research methods foundations (2019).
- Parra and Brusilovsky (2015) Denis Parra and Peter Brusilovsky. 2015. User-controllable personalization: A case study with SetFusion. International Journal of Human-Computer Studies 78 (2015), 43–67.
- Paymans et al. (2004) Tim F Paymans, Jasper Lindenberg, and Mark Neerincx. 2004. Usability trade-offs for adaptive user interfaces: ease of use and learnability. In IUI2004. 301–303.
- Peissner and Edlin-White (2013) Matthias Peissner and Rob Edlin-White. 2013. User control in adaptive user interfaces for accessibility. In Human-Computer Interaction–INTERACT 2013: 14th IFIP TC 13 International Conference, Cape Town, South Africa, September 2-6, 2013, Proceedings, Part I 14. Springer, 623–640.
- Persson et al. (2001) Anders Persson, Birgitta Wanek, and Anja Johansson. 2001. Passive versus active operator work in automated process control—a job design case study in a control centre. Applied ergonomics 32, 5 (2001), 441–451.
- Philip et al. (2022) Ben Joseph Philip, Mohamed Abdelrazek, Alessio Bonti, Scott Barnett, and John Grundy. 2022. Data collection mechanisms in health and wellness apps: review and analysis. JMIR mHealth and uHealth 10, 3 (2022), e30468.
- Pizzoli et al. (2019) Silvia Maria Francesca Pizzoli, Chiara Renzi, Paola Arnaboldi, William Russell-Edu, and Gabriella Pravettoni. 2019. From life-threatening to chronic disease: Is this the case of cancers? A systematic review. Cogent Psychology 6 (2019). https://api.semanticscholar.org/CorpusID:86785333
- Prasad et al. (2014) Aarathi Prasad, Jacob Sorber, Timothy Stablein, D Anthony, and David Kotz. 2014. Understanding user privacy preferences for mhealth data sharing. mHealth multidisciplinary verticals 30 (2014), 545–569.
- Pu and Chen (2006) Pearl Pu and Li Chen. 2006. Trust building with explanation interfaces. In Proceedings of the 11th international conference on Intelligent user interfaces. 93–100.
- Quiñones et al. (2018) Daniela Quiñones, Cristian Rusu, and Virginica Rusu. 2018. A methodology to develop usability/user experience heuristics. Computer standards & interfaces 59 (2018), 109–129.
- Ramos et al. (2019) Gonzalo Ramos, Jina Suh, Soroush Ghorashi, Christopher Meek, Richard Banks, Saleema Amershi, Rebecca Fiebrink, Alison Smith-Renner, and Gagan Bansal. 2019. Emerging perspectives in human-centered machine learning. In Extended Abstracts of the 2019 CHI Conference on Human Factors in Computing Systems. 1–8.
- Reinecke and Bernstein (2013) Katharina Reinecke and Abraham Bernstein. 2013. Knowing what a user likes: A design science approach to interfaces that automatically adapt to culture. Mis Quarterly (2013), 427–453.
- Rich and McGee (2004) Aaron Rich and Mick McGee. 2004. Expected usability magnitude estimation. In Human Factors and Ergonomics Society Annual Meeting, Vol. 48. SAGE Publications Sage CA: Los Angeles, CA, 912–916.
- Rivera (2005) Diego Rivera. 2005. The effect of content customization on learnability and perceived workload. In CHI2005. 1749–1752.
- Robbins et al. (2017) Rebecca Robbins, Paul Krebs, Ram Jagannathan, Girardin Jean-Louis, Dustin T Duncan, et al. 2017. Health app use among US mobile phone users: analysis of trends by chronic disease status. JMIR mHealth and uHealth 5, 12 (2017), e7832.
- Rosman et al. (2014) Benjamin Rosman, Subramanian Ramamoorthy, M. M. Hassan Mahmud, and Pushmeet Kohli. 2014. On user behaviour adaptation under interface change. IUI2014 (2014). https://api.semanticscholar.org/CorpusID:1724637
- Sanders and Martin-Hammond (2019) Jamie Sanders and Aqueasha Martin-Hammond. 2019. Exploring autonomy in the design of an intelligent health assistant for older adults. In Proceedings of the 24th International conference on intelligent user interfaces: companion. 95–96.
- Sara (2021) Berg Sara. 2021. What doctors wish patients knew about decision fatigue. https://www.ama-assn.org/delivering-care/public-health/what-doctors-wish-patients-knew-about-decision-fatigue
- Scherer et al. (2021) Laura D Scherer, Daniel D Matlock, Larry A Allen, Chris E Knoepke, Colleen K McIlvennan, Monica D Fitzgerald, Vinay Kini, Channing E Tate, Grace Lin, and Hillary D Lum. 2021. Patient roadmaps for chronic illness: Introducing a new approach for fostering patient-centered care. MDM Policy & Practice 6, 1 (2021), 23814683211019947.
- Schiavo et al. (2022) Gianluca Schiavo, Ornella Mich, Michela Ferron, and Nadia Mana. 2022. Trade-offs in the design of multimodal interaction for older adults. Behaviour & Information Technology 41, 5 (2022), 1035–1051.
- Schmid-Isler (2000) Salome Schmid-Isler. 2000. The language of digital genres-a semiotic investigation of style and iconology on the world wide web. In Proceedings of the 33rd Annual Hawaii International Conference on System Sciences. IEEE, 9–pp.
- Setiawan et al. (2019) I Made Agus Setiawan, Leming Zhou, Zakiy Alfikri, Andi Saptono, Andrea D Fairman, Brad Edward Dicianno, and Bambang Parmanto. 2019. An adaptive mobile health system to support self-management for persons with chronic conditions and disabilities: usability and feasibility studies. JMIR Formative Research 3, 2 (2019), e12982.
- Shakshuki et al. (2015) Elhadi M Shakshuki, Malcolm Reid, and Tarek R Sheltami. 2015. An adaptive user interface in healthcare. Procedia Computer Science 56 (2015), 49–58.
- Short et al. (2018) Camille E Short, Ann DeSmet, Catherine Woods, Susan L Williams, Carol Maher, Anouk Middelweerd, Andre Matthias Müller, Petra A Wark, Corneel Vandelanotte, Louise Poppe, et al. 2018. Measuring engagement in eHealth and mHealth behavior change interventions: viewpoint of methodologies. Journal of medical Internet research 20, 11 (2018), e292.
- Siala et al. (2004) Haytham Siala, Robert M O’Keefe, and Kate S Hone. 2004. The impact of religious affiliation on trust in the context of electronic commerce. Interacting with computers 16, 1 (2004), 7–27.
- Spann and Stewart (2018) Alice Spann and Ellen Stewart. 2018. Barriers and facilitators of older people’s mHealth usage: a qualitative review of older people’s views. Human Technology: An Interdisciplinary Journal on Humans in ICT Environments 14, 3 (2018), 264–296.
- Sproull et al. (1996) Lee Sproull, Mani Subramani, Sara Kiesler, Janet H Walker, and Keith Waters. 1996. When the interface is a face. Human-computer interaction 11, 2 (1996), 97–124.
- Srivastava et al. (2021) Ayushi Srivastava, Shivani Kapania, Anupriya Tuli, and Pushpendra Singh. 2021. Actionable UI Design Guidelines for Smartphone Applications Inclusive of Low-Literate Users. Proceedings of the ACM on Human-Computer Interaction 5, CSCW1 (2021), 1–30.
- Stawarz et al. (2014) Katarzyna Stawarz, Anna L Cox, and Ann Blandford. 2014. Don’t forget your pill! Designing effective medication reminder apps that support users’ daily routines. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. 2269–2278.
- Stephan et al. (2017) Laura Siga Stephan, Eduardo Dytz Almeida, Raphael Boesche Guimaraes, Antonio Gaudie Ley, Rodrigo Gonçalves Mathias, Maria Valéria Assis, and Tiago Luiz Luz Leiria. 2017. Processes and recommendations for creating mHealth apps for low-income populations. JMIR mHealth and uHealth 5, 4 (2017), e6510.
- Strauss and Corbin (1994) Anselm Strauss and Juliet Corbin. 1994. Grounded theory methodology: An overview. (1994).
- Sun et al. (2013) Yongqiang Sun, Nan Wang, Xitong Guo, and Zeyu Peng. 2013. Understanding the acceptance of mobile health services: a comparison and integration of alternative models. Journal of electronic commerce research 14, 2 (2013), 183.
- Sundar (2008) S Shyam Sundar. 2008. Self as source: Agency and customization in interactive media. In Mediated interpersonal communication. Routledge, 72–88.
- Sundar and Marathe (2010) S Shyam Sundar and Sampada S Marathe. 2010. Personalization versus customization: The importance of agency, privacy, and power usage. Human communication research 36, 3 (2010), 298–322.
- Talboom-Kamp et al. (2018) Esther PWA Talboom-Kamp, Noortje A Verdijk, Marise J Kasteleyn, Mattijs E Numans, and Niels H Chavannes. 2018. From chronic disease management to person-centered eHealth; a review on the necessity for blended care. Clinical eHealth 1, 1 (2018), 3–7.
- Taylor and Todd (1995) Shirley Taylor and Peter Todd. 1995. Assessing IT usage: The role of prior experience. MIS quarterly (1995), 561–570.
- Tinschert et al. (2017) Peter Tinschert, Robert Jakob, Filipe Barata, Jan-Niklas Kramer, Tobias Kowatsch, et al. 2017. The potential of mobile apps for improving asthma self-management: a review of publicly available and well-adopted asthma apps. JMIR mHealth and uHealth 5, 8 (2017), e7177.
- Torous et al. (2015) John Torous, Patrick Staples, and Jukka-Pekka Onnela. 2015. Realizing the potential of mobile mental health: new methods for new data in psychiatry. Current psychiatry reports 17 (2015), 1–7.
- Trajkova and Martin-Hammond (2020) Milka Trajkova and Aqueasha Martin-Hammond. 2020. ” Alexa is a Toy”: exploring older adults’ reasons for using, limiting, and abandoning echo. In CHI2020. 1–13.
- Tsandilas and Schraefel (2005) Theophanis Tsandilas and Monica M. C. Schraefel. 2005. An empirical assessment of adaptation techniques. CHI2005 (2005). https://api.semanticscholar.org/CorpusID:18812760
- Turk et al. (2019) Amadea Turk, Emma Fairclough, Gillian Grason Smith, Benjamin Lond, Veronica Nanton, and Jeremy Dale. 2019. Exploring the perceived usefulness and ease of use of a personalized web-based resource (care companion) to support informal caring: qualitative descriptive study. JMIR aging 2, 2 (2019), e13875.
- Vajawat et al. (2021) Bhavika Vajawat, Prateek Varshney, and Debanjan Banerjee. 2021. Digital gaming interventions in psychiatry: evidence, applications and challenges. Psychiatry research 295 (2021), 113585.
- Varshney (2014) Upkar Varshney. 2014. Mobile health: Four emerging themes of research. Decision Support Systems 66 (2014), 20–35.
- Vasilyeva et al. (2005) Ekaterina Vasilyeva, Mykola Pechenizkiy, and Seppo Puuronen. 2005. Towards the framework of adaptive user interfaces for eHealth. In 18th IEEE Symposium on Computer-Based Medical Systems (CBMS’05). IEEE, 139–144.
- Venkatesh et al. (2012) Viswanath Venkatesh, James YL Thong, and Xin Xu. 2012. Consumer acceptance and use of information technology: extending the unified theory of acceptance and use of technology. MIS quarterly (2012), 157–178.
- Vijan (2015) Sandeep Vijan. 2015. Type 2 diabetes. Annals of Internal Medicine 162, 5 (2015), ITC1–ITC16.
- Vogt and Meier (2010) Joël Vogt and Andreas Meier. 2010. An adaptive user interface framework for eHealth services based on UIML. (2010).
- Wang et al. (2023a) Wei Wang, Hourieh Khalajzadeh, John Grundy, and Anuradha Madugalla. 2023a. Adaptive user interfaces for software supporting chronic diseases. In 2023 IEEE Symposium on Visual Languages and Human-Centric Computing (VL/HCC). IEEE, 215–218.
- Wang et al. (2023b) Wei Wang, Hourieh Khalajzadeh, John Grundy, Anuradha Madugalla, Jennifer McIntosh, and Humphrey O Obie. 2023b. Adaptive user interfaces in systems targeting chronic disease: a systematic literature review. User Modeling and User-Adapted Interaction (2023), 1–68.
- Wang et al. (2024) Wei Wang, Hourieh Khalajzadeh, John Grundy, Anuradha Madugalla, and Humphrey Obie. 2024. Adaptive User Interfaces for Software Supporting Chronic Disease. In 2024 IEEE/ACM 45th International Conference on Software Engineering: Software Engineering in Society (ICSE-SEIS). IEEE.
- Wesson et al. (2010) Janet L Wesson, Akash Singh, and Bradley Van Tonder. 2010. Can adaptive interfaces improve the usability of mobile applications?. In Human-Computer Interaction: Second IFIP TC 13 Symposium, HCIS 2010, Held as Part of WCC 2010, Brisbane, Australia, September 20-23, 2010. Proceedings. Springer, 187–198.
- WHO (2022) WHO. 2022. Invisible numbers: the true extent of noncommunicable diseases and what to do about them. Retrieved September 22, 2022 from https://www.who.int/publications/i/item/9789240057661
- WHO (2023) World Health Organization: WHO. 2023. Noncommunicable diseases. https://www.who.int/news-room/fact-sheets/detail/noncommunicable-diseases
- Wilson (2014) Chauncey Wilson. 2014. Chapter 2 - Semi-Structured Interviews. In Interview Techniques for UX Practitioners, Chauncey Wilson (Ed.). Morgan Kaufmann, Boston, 23–41. https://doi.org/10.1016/B978-0-12-410393-1.00002-8
- Woldaregay et al. (2018) Ashenafi Zebene Woldaregay, David-Zacharie Issom, André Henriksen, Henna Marttila, Martin Mikalsen, Gerit Pfuhl, Keiichi Sato, Christian Lovis, and Gunnar Hartvigsen. 2018. Motivational Factors for User Engagement with mHealth Apps.. In pHealth. 151–157.
- Xie (2009) Bo Xie. 2009. Older adults’ health information wants in the internet age: implications for patient–provider relationships. Journal of health communication 14, 6 (2009), 510–524.
- Xu et al. (2014) Jie Xu, Kim Le, Annika Deitermann, and Enid Montague. 2014. How different types of users develop trust in technology: A qualitative analysis of the antecedents of active and passive user trust in a shared technology. Applied ergonomics 45, 6 (2014), 1495–1503.
- Yuan and Herbert (2012) Bingchuan Yuan and John Herbert. 2012. A fuzzy-based context modeling and reasoning framework for CARA pervasive healthcare. In Impact Analysis of Solutions for Chronic Disease Prevention and Management: 10th International Conference on Smart Homes and Health Telematics, ICOST 2012, Artiminio, Italy, June 12-15, 2012. Proceedings 10. Springer, 254–257.
- Zaphiris et al. (2007) Panayiotis Zaphiris, Sri Kurniawan, and Mariya Ghiawadwala. 2007. A systematic approach to the development of research-based web design guidelines for older people. Universal Access in the information Society 6 (2007), 59–75.
- Zaphiris et al. (2009) Panayiotis Zaphiris, Ulrike Pfeil, and Dorian Xhixho. 2009. User evaluation of age-centred web design guidelines. In Universal Access in Human-Computer Interaction. Addressing Diversity: 5th International Conference, UAHCI 2009, Held as Part of HCI International 2009, San Diego, CA, USA, July 19-24, 2009. Proceedings, Part I 5. Springer, 677–686.
- Zeidler et al. (2013) Clemens Zeidler, Christof Lutteroth, and Gerald Weber. 2013. An evaluation of advanced user interface customization. In OZCHI2013. 295–304.
- Zhang et al. (2021) Renwen Zhang, Kathryn E. Ringland, Melina Paan, David C. Mohr, and Madhu Reddy. 2021. Designing for emotional well-being: integrating persuasion and customization into mental health technologies. In CHI2021. 1–13.
Appendix A Interview and focus group post-session questions
Introduction 2023 Interview/focus group post-session survey Thanks for attending today’s interview/focus group session. We invite you to complete a post-session survey to further share your thoughts on the adaptive user interface for chronic disease-related applications.
Please enter your withdrawal code below (You can
To what extent do you want to have the following main types of adaptation in your daily use of mHealth applications? (Much weaker, Moderately weaker,Slightly weaker, Neutral, Slightly stronger, Moderately stronger, Much stronger)
Presentation adaptation
Content adaptation
Behaviour adaptation
To what extent do you want to have the following presentation adaptation in your daily use of mHealth applications? (Much weaker, Moderately weaker,Slightly weaker, Neutral, Slightly stronger, Moderately stronger, Much stronger)
Graphic design
Color and theme
Font size
Avatar/animation
Information architecture
To what extent do you want to have the following content adaptations in your daily use of mHealth applications (Much weaker, Moderately weaker,Slightly weaker, Neutral, Slightly stronger, Moderately stronger, Much stronger)?
Interface elements rearrangement
Reorder the elements
Hide the elements
Content complexity
Easy to understand language
Minmalize the design
Change text to image
To what extent do you want to have the following behaviour adaptations in your daily use of mHealth applications (Much weaker, Moderately weaker,Slightly weaker, Neutral, Slightly stronger, Moderately stronger, Much stronger)?
Navigation adaptation
Add on functions
Different persuasive strategy
Difficulty levels
Multimodal interaction
Which kind of data do you prefer to use for your interface adaptation?
Your role
Your feedback
Your motivation
Your interaction with the interface
Your game performance
Your physical characteristics
Your psychological characteristics
Your demographic information
Your preference
Your emotions
Your social activity
Your physiological characteristics
your goals
How do you wish your data to be collected?
Analyse of user behaviour through the application
Analysis of activities with keyboard
Smartphone sensor
User input through the application
External sensor
Wearable sensor
What level of initiative do you want to take during the adaptation you prefer?
Manual system
Automatic system
Semi-automatic system
Appendix B Survey questions
Introduction AUIs are software applications where some aspect(s) of the interface is modified to cater for different user needs and/or preferences e.g. font or button size, colour, layout, complexity, interaction style and so on. We are researching how AUIs can be leveraged to better cater for users with chronic diseases. As part of this survey, you will be asked to answer questions about your perspective toward AUI. On average it will take 10-15 minutes for a participant to complete the survey. You can enter a draw for a AU$20 voucher if you complete an online survey. Your email address will not be associated with your response since the contact information is collected in another survey. At the end of the survey, you will also be asked to indicate whether you are interested in participating in a focus group study. You may register through this link if you are interested in the focus group study. If you have any questions, please email wei.wang5@monash.edu. Ethical approval has been provided by Monash University. Further details can be provided upon request.
Survey Consent Form
I have read and understood the Explanatory Statement and I hereby consent to answer the online survey.
I understand that all information I provide for this study will be treated confidentially.
I understand that I am free to contact any of the people involved in the research to seek further clarification and information.
I agree that the data I provide in this research may be used by the investigators in future research projects.
Section 1:Demographic questions. This section will collect your demographic information. We will not share any identifying information that you submit. The demographic information collected is used only to assess the representatives of the survey participants.
In what age group are you?
18 - 24
25 - 34
35 - 44
45 - 54
55 - 64
65 - 74
75 - 84
85 or older
To which gender identity do you most identify as?
Male
Female
Non-binary / third gender
Prefer not to say
Other_
In which country do you currently reside?
China … Zimbabwe
What is your highest educational qualification?
Some high school credit, no diploma, or the equivalent
High school graduate, diploma, or the equivalent
Some college credit, no degree
Trade/technical/vocational training
Associate degree
Bachelor’s degree
Master’s degree
Doctorate degree
Other_
Section 2: Your health status. This section will collect your health information.
Has a health care provider ever told you have a chronic disease?
I have chronic disease
Other_
What chronic disease do you have?_
Section 3: mHealth applications. This section will collect information for your usage for mHealth applications.
Have you used mHealth apps before to manage your chronic disease? (The most common application of mHealth is the use of mobile devices to educate consumers about preventive healthcare services. However, mHealth is also used for disease surveillance, treatment support, epidemic outbreak tracking and chronic disease management.)
Yes, I have used/curently use the mHealth application
No, I never used before
Other _
What kind of mHealth application(s) have you used before?
Assistive application help users complete a task or activity more easily
Communication application to communicate with carers, family members, doctors, or emergency contacts
Informative application
Health game application (games that encourage exercise, good living habits)
Health promoting and self-monitoring application
Rehabilitation application
Other (Please specify the name of the application)
Why do you mainly use these application(s) for?
Health promoting and self-monitoring application
Rehabilitation application
Other (Please specify the name of the application)
To track what I eat
Help with weight loss
Manage my medications
Increase my physical activity levels
To monitor my chronic disease symptoms
To get education about my chronic disease
Other (Please specify)
How frequently do you use health apps?
less than once a month
Once a week
Less than once a week
2 or more times per day
How long each time you use the app?
1-10 minutes
10-20 minutes
more than 30 minutes
more than 60 minutes
Section 4:Adaptive user interface. This section will collect your perspective on the AUI. An AUI is a user interface which adapts its layout and elements to the needs of the user or context. Here are two examples of AUIs:
Figure 13. Exercise application with AUI
Figure 14. PD-Helper for Parkinson’s Disease Patients Exercise application (Figure 13): The left side is a standard version of the interface, to maintain usability and readability of the interface when the user is running, the interface automatically hide images and changes small-font and enlarge the subtitle and two buttons. (park2015model).
PD-Helper for Parkinson’s Disease Patients (Figure 14): Users are allowed to change the font size and other settings in accordance with their own preferences and capabilities (Jabeen et al., 2019).-
Which kind of adaptation do you prefer for interface adaptation? (mHealth apps can be adapted based on your needs in different ways. Select all that apply)
Graphic design (layout, font size, colours and themes)
Navigation adaptation (only specified functions can be used to provide help in special cases)
Different persuasive strategy (to better motivate the desired behaviour change, according to different user types)
Content complexity (making content easy to understand and process)
Information architecture (giving the user more freedom when navigating through large amounts of textual information)
Multimodal interaction (voice/text input, switching of voice/text output in different environments)
Interface elements rearrangement (removing, adding or reordering elements on the page)
Difficulty level (change difficulty level of game/exercise based on user’s motivation or performance)
Add on functions (to better assist the user in using the application, e.g. zoom function for older users)
Sound effect (adjust the volume depending on the distance of the person from the device)
Other (Please specify)
Which kind of data do you prefer to use for your interface adaptation? (Imagine your mHealth app can be adapted based on your characteristics and preference. Please see different types of data that can be used in adapting mHealth apps. Select all that apply)
Your role (you will perform different jobs and be in different situations
Your feedback (your likes, dislikes and preferences for the interface)
Your motivation (what motivates you to use the application)
Your interaction with the interface (e.g. number of clicks, links visited, time spent browsing, etc.)
Your game performance (scores, successes and wins in the game)
Your physical characteristics (ability to perform different activities in daily life)
Your psychological characteristics (thoughts, personality, feelings and other cognitive characteristics)
Your demographic information (quantifiable insights of users into the population such as age, gender and education level)
Your preferences (preferred layout, input/output, theme and interface design)
Your emotions (based on emotions when using the app)
Your social activity (the extent to which you interact with others around you)
Your physiological characteristics (e.g. stress levels, heart rate, blood pressure and blood oxygen levels)
Your goals (the end state you want to achieve by using the app)
Other (Please specify)
How do you wish your data to be collected? (Select all that apply)
Analysis of user behaviour through the application (checking the history of phone usage, game performance)
Analysis of activities with keyboard
Smartphone sensor (phone camera, accelerometer, GPS, microphone)
User input through the application (user manually input their interest, gender, preference)
External sensor (Kinect sensor, computer/TV camera)
Wearable sensor (pedometer, smartwatch/bracelet, medical sensors (blood pressure monitor…)
Other (Please specify)
What level of initiative do you want to take during the adaptation you are preferred?
Manual system (allows you to manually modify certain settings of the user interface)
Automatic system (the system adjusts the user interface automatically)
Semi-automatic system (you and the system collaborate to achieve adaptation)
Other (Please specify)
How did you hear about us?
Linkedin post
Facebook post/group
Twitter post
Email
Word of mouth
Baker Institute Diabetes Clinic
Alfred Hospital
Dementia Australia, Stroke Foundation, or Kidney Australia
Other social media_
Is there anything else you want to tell us about this survey or our research study?_
Would you like us to email you the survey result? (if yes, please leave your email below)